### v1

# Generalized Label-Efficient 3D Scene Parsing via Hierarchical Feature Aligned Pre-Training and Region-Aware Fine-tuning

Abstract
Deep neural network models have achieved remarkable progress in 3D scene understanding while trained in the closed-set setting and with full labels. However, the major bottleneck for the current 3D recognition approach is that these models do not have the capacity to recognize any unseen novel classes beyond the training categories in diverse kinds of real-world applications. In the meantime, current state-of-the-art 3D scene understanding approaches primarily require a large number of high-quality labels to train neural networks, which merely perform well in a fully supervised manner. Therefore, we are in urgent need of a framework that can simultaneously be applicable to both 3D point cloud segmentation and detection, particularly in the circumstances where the labels are rather scarce. This work presents a generalized and simple framework for dealing with 3D scene understanding when the labeled scenes are quite limited. To extract knowledge for novel categories from the pre-trained vision-language models, we propose a hierarchical feature-aligned pre-training and knowledge distillation strategy to extract and distill meaningful information from large-scale vision-language models, which helps benefit the open-vocabulary scene understanding tasks. To leverage the boundary information, we propose a novel energy-based loss with boundary awareness benefiting from the region-level boundary predictions. To encourage latent instance discrimination and to guarantee efficiency, we propose the unsupervised region-level semantic contrastive learning scheme for point clouds, using confident predictions of the neural network to discriminate the intermediate feature embeddings at multiple stages. In the limited reconstruction case, our proposed approach, termed WS3D++, ranks 1st on the large-scale ScanNet benchmark on both the task of semantic segmentation and instance segmentation. Also, our proposed WS3D++ achieves state-of-the-art data-efficient learning performance on the other large-scale real-scene indoor and outdoor datasets S3DIS and SemanticKITTI. Extensive experiments with both indoor and outdoor scenes demonstrated the effectiveness of our approach in both data-efficient learning and open-world few-shot learning. All codes, models, and data are to made publicly available at: https://github.com/KangchengLiu. The code is at: WS3D++ Code link.
Index Terms:
3D Scene Understanding, Data-efficient Learning, Region-Level Contrast, Energy Function, 3D Vision-Language Model
1 Introduction
The 3D scene parsing problem, which typically encompasses several important downstream tasks: point cloud semantic segmentation, instance segmentation, and object detection, becomes increasingly important with the wide deployment of 3D sensors, such as LiDAR and RGB-D cameras [1, 2, 3, 4, 5, 6, 7, 8]. Point clouds are raw sensor data obtained from 3D sensors and the most simple and common 3D data representation for understanding 3D scenes of robot navigation, robot grasping, and manipulation tasks. Despite significant success in deep neural networks applied to 3D visual perception, two major challenges hinder the construction of more scalable visual perception systems in 3D worlds. One is the closed-set assumption, which means the model only performs well while recognizing the categories that appear in the training set and struggles in recognizing the novel unseen object categories or concepts. Another is the heavy reliance on large amounts of high-quality labeled data. Large-scale 3D scenes are very laborious to label, which also makes it very hard for deep network models to perform well with very limited annotations.
Close-set assumption: One of the major bottlenecks in scaling up visual perception systems is the poor generalization capacity while encountered with diverse novel semantic classes or severe domain shifts. To endow the model with the capacity for adapting the learned representation and make it conform to different data distributions as well as recognize diverse novel categories, pioneer researches such as CLIP [9], Flamingo [10], and Otter [11] have demonstrated the great potentials in learning well-aligned visual linguistic representation from large-scale image-text pairs on the Internet for improving the model generalization capacity. To this end, subsequent approaches have been proposed in establishing abundant vision-language associations for different visual recognition tasks including detection and segmentation using the large-scale vision-language model (VLM) [12, 13, 14]. The paired visual-linguistic feature representation can enable the recognition of a large number of novel objects or concepts with natural language supervision because the visual and the lexical language features are well-matched in their shared semantic feature space. Despite the remarkable performance achieved in developing diverse vision-language foundation models such as SAM [15] and SEEM [16] for image-based scene understanding, it remains very difficult for CLIP [9] to benefit downstream 3D scene understanding because it is difficult to raise the feature dimension to 3D and establish explicit correlations or find clear alignments between large-scale scene/object-level 3D point clouds as well as linguistic semantic concepts. Moreover, it is even harder to transfer the informative knowledge to various downstream 3D scene understanding tasks. These limitations severely restrict the scalability of VLM to handle diverse unseen 3D scenes containing diverse novel 3D object categories.
Reliance on large-scale labeled data. A critical prerequisite for fully exploiting the capacity of the fully supervised deep learning approaches is the accessibility to large-scale well-annotated high-quality training data. Most point cloud understanding methods rely on heavy annotations [17, 18, 19]. However, the annotation of large-scale 3D point cloud scenes is rather time-consuming and labor-intensive. For instance, it requires around thirty minutes to label a single scene for ScanNet [20] or S3DIS [21] with thousands of scenes. Though existing point cloud understanding methods [17, 18, 19] have achieved decent results on these datasets, it is difficult to directly extend them to novel scenes when the high-quality labeled data is scarce. In the meanwhile, it is often the case that a limited number of scenes can be reconstructed in real applications [22]. Therefore, developing methods that can be trained with very limited labeled scenes, termed data-efficient 3D scene understanding with limited scene-level annotation, becomes in high demand. Data-efficient semantic and instance segmentation [23, 24, 25] is a vehement research topic for image-level scene understanding. Some simple but successful methods have been proposed, such as contrastive learning [26, 27] which learns a meaningful and discriminative representation, and conditional random field (CRF) [28, 29] for pseudo label propagation. However, there still exist four main challenging unsolved issues while scaling up these approaches to 3D scene understanding. First, the widely adopted energy function-based conditional random field segmentation [29] relies on handcrafted feature similarities and does not consider explicit boundary information. It attaches equal importance to pixels on semantic boundaries and within same semantic objects, which can cause vague and inaccurate predictions in pixel-level segmentation at object boundaries. And how to leverage boundary information has been explored in 2D but rarely explored in 3D data efficient learning [30]. Second, the computation costs are both very high when applying point-level contrastive learning or point-level energy-based segmentation in a dense point cloud scene for every point pair [31]. Furthermore, large-scale point cloud scenes even contain billions of points, making point-level contrastive learning intractable in computational costs. Third, the existing unsupervised contrastive learning-based pre-training for point clouds [27, 22, 32, 33] only considers geometrically registered point/voxel pairs as the positive samples, while it does not explicitly consider explicit regional information, let alone the hierarchical alignments.
Figure 1: Overall illustrative diagram of our proposed WS3D++. We integrate language-3D feature associated pre-training and data-efficient fine-tuning to achieve effective data-efficient and open-vocabulary 3D scene understanding for 3D scenes.
Driven by the above motivations in terms of both generalization capacity and data efficiency, we propose an effective two-stage framework, involving unsupervised hierarchical vision-language pre-training and label-efficient fine-tuning to enable more label/data-efficient 3D scene understanding. As shown in Fig. 1, in the pre-training stage, we leverage the rendering techniques to construct well-aligned 2D views for large-scale 3D scenes to establish more accurate coarse-to-fine vision-language associations. Then, we leverage the off-the-shelf object detector and the pre-trained large-scale vision-language model CLIP [9] to construct the hierarchical feature representations from both the global scene level to the local object level. We also propose an effective knowledge distillation strategy that distills the informative visual-language-aligned representation of the image encoder in CLIP [9] to the 3D backbone network.
During the fine-tuning stage, as shown in Fig. 1, we propose a unified WS3D++ framework that simultaneously solves the 3D scene understanding problem under the data-efficient setting. We first use the over-segmentation [34] to obtain regions and use a boundary prediction network (BPN) as an intermediate tool to obtain boundary region labels. Then, high-confidence boundary region labels serve as the guidance for our proposed region-level energy-based loss. Meanwhile, we propose a region-level confidence-guided contrastive loss to enhance instance discrimination. Specifically, our WS3D++ includes two innovative designs to address the very challenging label scarcity issues and to enhance performance. Firstly, to encourage latent instance discrimination and to guarantee efficiency, an efficient region-level feature contrastive learning strategy is proposed to guide network training at multiple stages, which realizes the unsupervised instance discrimination. Also, to leverage boundary information as labels for the final semantic divisions, an energy-based loss with guidance from the semantic boundary regions is proposed to take the maximized advantage of the unlabeled data in network training. Combined with supervised loss, the labeled data can also be leveraged to boost the final downstream 3D scene understanding performance.
WS3D++ is a significant extension of the preliminary version of the conference work WS3D [4], where basic ideas of boundary awareness and contrastive instance discrimination are introduced to tackle the data-efficient 3D scene understanding during fine-tuning stage. In summary, we extensively enriched previous works in the following aspects:
First, we propose a generalized pre-training approach for data-efficient learning, which establishes accurate alignments between language and 3D point cloud in terms of both object-level and scene-level semantics in a hierarchical manner.  Second, we propose to leverage rendering techniques which make explicit associations between image and point cloud to facilitate 2D-to-3D matching and subsequent language-to-3D matching.  Third, we visualized the language-queried activation maps directly on the 3D scenes, which demonstrates that the proposed approach learns better visual-linguistic alignment between the language descriptions and the visual object-level information. Finally, we evaluate our proposed approach comprehensively in diverse data-efficient and open-world learning settings for both the 3D semantic segmentation and 3D instance segmentation tasks.
The contributions of our work are highlighted as follows:
  1. (1)
    During the pre-training stage, we first design an effective approach which distills rich knowledge from the large-scale vision-language model into the 3D point cloud modality. Specifically, we propose leveraging rendering to obtain explicit scene-level and object-level 2D-3D feature associations, establishing a more accurate vision-language association hierarchically than the original CLIP encoder.
  2. (2)
    During the pre-training stage, we further propose a word-to-3D matching approach to establish the well-aligned language-3D associated feature representations at both the scene level and the object level, which largely facilitates the subsequent effective contrastive learning with the mostly matched visual-linguistic contrastive pairs. The proposed designs have both enhanced the data-efficient learning and the knowledge-transfer capacity of the model, as demonstrated by our extensive experiments on both the 3D object detection and the 3D semantic/instance segmentation tasks.
  3. (3)
    During the fine-tuning stage, we propose a region-aware energy-based optimization approach to achieve the region-level boundary awareness, which utilizes the boundary as additional information to help assist the 3D scene segmentation and understanding. Furthermore, we propose the unsupervised region-level semantic contrastive learning strategy for multi-stage feature discriminations. The energy-based loss and the contrastive loss are jointly optimized for pre-training the backbone network in a complementary manner, which take full advantage of the unlabeled data.
  4. (4)
    Integrating the above two stages as a whole, we propose a unified framework termed WS3D++. State-of-the-art performance has been achieved by it with extensive experiments conducted on ScanNet [20] and other indoor/outdoor benchmarks such as S3DIS [21], SemanticKITTI [18] and NuScenes [35] in diverse experimental settings without bells and whistles. Finally, our proposed approach achieves pioneer performance on the very large-scale ScanNet [20] dataset in diverse downstream tasks of 3D scene understanding, including tasks among 3D semantic segmentation, 3D instance segmentation, and 3D object detection [^1] http://kaldir.vc.in.tum.de/scannet_benchmark/data_efficient/.
To the best of our knowledge, this is the pioneer work which comprehensively evaluates across diverse 3D label-efficient scene understanding downstream tasks with our proposed 3D open-vocabulary recognition approach termed WS3D++. Our endeavor is orthogonal to the 3D backbone network designs and thus can be seamlessly integrated with the prevailing 3D point cloud detection or segmentation models. Our comprehensive results provide solid baselines for future research in data-efficient 3D scene understanding.
2 Related Work
Learning-based Point Cloud Understanding. Deep-network-based approaches are widely adopted for point cloud understanding. Fully supervised approaches can be roughly categorized into voxel-based [36, 37, 38, 39, 40], projection-based [41, 42, 43, 44, 45, 46], and point-based approaches [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. The voxel-based approaches [59, 60, 61] which are built upon SparseConv [59] and voxelize the point cloud for efficient processing have achieved remarkable performance in 3D scene parsing. Therefore, we use SparseConv [59] as our backbone architecture for downstream semantic understanding tasks because of its high performance in inferring 3D semantics. Also, the LiDAR-based approaches are of significance to many industrial applications such as inspections [62, 63, 64] and robot localization [65, 66, 67, 68, 69], and large-scale robotic scene parsing [57, 70, 71], as well as robotic control applications [72, 73, 74, 75, 76, 77], etc.
Pre-training for 3D Representation Learning. Many recent works propose to pre-train networks on source datasets with auxiliary tasks such as low-level point cloud geometric registration [27], 3D local structural prediction [78], the completion of the occluded point clouds [79], and the foreground-background feature discrimination [58], with effective learning strategies such as contrastive learning [27] and masked generative modelling [80, 81]. Then they finetune the weights of the trained networks for the downstream target tasks to boost their performances. However, several major challenges still exist. First, the above pre-training approaches all rely on the closed-set assumption, which means that the model can barely be transferred to recognize novel categories that do not appear within the training data. Second, the above methods require accessibility to the well-registered augmented point cloud [27, 22, 32, 33] to construct the pre-training contrast views, which are very hard to obtain for large-scale 3D scenes. Third, a large number of computational power is required in the pre-training stage. Therefore, the designed pre-training approaches need to be very simple and lightweight, thus will make it easier to directly transfer the model to large-scale point clouds.
Recently, with the development of large-scale vision-language models such as CLIP [9] and Flamingo [10], we can largely benefit the recognition capacity of 3D scene understanding models by distilling knowledge from large-scale vision-language models. For example, some pioneer works such as PointCLIP [82, 83] and ULIP [84] successfully transfer the knowledge from vision-language models to boost the downstream 3D shape classification tasks. The 3D CAD shapes are transferred to multi-view depth maps, thus they can be fed into the CLIP visual encoder and used the image representations paired with the corresponding point cloud as a bridge to obtain the correlations between the 3D and textual features. Moving beyond object-level recognition tasks, pioneer works also explore how to establish alignment among images, language, and 3D point cloud scenes for the task of open-vocabulary 3D scene understanding [85, 86, 87], which target localizing, detecting, and segmenting novel object categories that do not exist in the annotation. Compared with them, our proposed simple but effective framework can be both applicable to data-efficient learning and open-vocabulary scene understanding.
Figure 2: The pre-training paradigm of our proposed WS3D++. We propose hierarchical feature alignments to establish hierarchical vision-language aligned feature representations during pre-training. This kind of paradigm helps learn more powerful visual-linguistic matched representations ranging from both the global view-level to the local object category-level.
Label/Data-Efficient Learning for 3D. Recent studies have produced many elaborately designed backbone networks for 3D semantic/instance segmentation [88, 89, 45, 90, 91], as well as for 3D object detection [92, 93, 94, 95]. However, they rely on full supervision. Directly applying current SOTAs (State-of-the-art) methods for training will result in a great decrease in performance [96] for WSL, if the percentage of labeled data drops to a certain value, e.g., less than 30%. Recently, many works have started to focus on point cloud semantic segmentation with partially labeled data. Wang et al. [97] choose to transform point clouds to images, but pixel-level semantic segmentation labels are required for network training. Sub-cloud annotations [98] require extra labor to separate the sub-clouds and to label points within the sub-clouds. Liu [99] proposed a robust data-efficient 3D scene parsing framework. It leverages the complementary merits of the superior generalization capacity of the traditional 3D descriptors and the strong feature description capacity learned 3D descriptors to learn very robust local features. Then using the descriptor guided learned region merging [3], superior performance can be achieved on downstream tasks. Liu et al. [100, 101] proposed self-training techniques to tackle scene understanding in weak supervision, which design a two-stage training scheme to produce iteratively optimized pseudo labels from weak labels during training. Despite satisfactory results, these approaches still have not learned generalized representation applicable to diverse tasks. Xu et al. [102] adopt a weakly-supervised training strategy, which combines training with coarse-grained information and partial points using 10% labels. However, their tested cases are limited to object part segmentation, and it is difficult to uniformly choose points to label. The convex decomposition [103] is conducted in an approximate manner to perform 3D scene parsing on the object parts. More approaches [104] have been proposed recently, which utilize class prototypes and masked point cloud modeling [105, 106, 81] to learn informative representations for downstream 3D scene understanding. To sum up, although approaches have been proposed to alleviate the data efficiency problem, the models for weakly supervised learning lack the capacity to recognize novel categories beyond the labeled training set. Our framework tackles both open-set and data-efficient learning problems and is widely applicable to diverse 3D scene understanding downstream tasks.
3 Proposed Methodology
We propose a general WS3D++ framework to tackle weakly supervised 3D understanding with limited labels, as demonstrated in Fig. 1. Our framework consists of both the vision-language pre-training stage shown in Fig. 2 and fine-tuning stage illustrated in Fig. 3. During the pre-training stage, we first propose the hierarchical contrastive learning strategy with the help from rendering for more accurate vision-language alignments at both the scene level and object level. Then we also design a distillation strategy to distill point-language-aligned representations from 2D image network to 3D point-cloud network to endow 3D networks with the open-vocabulary recognition capacity. Finally, during the fine-tuning stage, we directly propose a weakly supervised approach, we perform fine-tuning with regional boundary awareness and region-aware instance discrimination, which significantly improve model discrimination capacities when the labeled data are rather scarce.
Figure 3: The fine-tuning paradigm of our proposed WS3D++, i.e. WS3D. WS3D [4] consists of three modules: 1. The unsupervised region-level energy-based optimization guided by boundary labels; 2. The unsupervised multi-stage region-level contrastive learning with high confidence; 3. The supervised region-level semantic contrastive learning with labeled data. The backbone network adopts encoder-decoder structures. The weights of the backbone network are shared in the supervised and unsupervised branches. Integrated with the proposed pre-training paradigm illustrated in Figure 2, by our proposed hierarchical feature aligned pre-training and regional fine-tuning, more effective label-efficient learning as well as open-vocabulary learning is realized.
3.1 Hierarchical Vision-Language Knowledge Associations and Distillations for Pre-training
We propose a hierarchical alignment strategy in pre-training, which employs the rendering approach as a bridge to effectively align 3D vision and language embeddings, thus capturing coarse-to-fine associations for visual-linguistic synergized representations from the global scene level to local object level. It enables extracting more accurate 3D-language associations in a hierarchical manner.
Multi-view rendering. To obtain paired 2D-3D representations, we propose leveraging multi-view rendering to obtain paired 2D views from 3D point cloud scenes. The pairing process consists of two steps, the first is to convert point cloud scenes into meshes and the second is to render 2D images based on the different views of the 3D meshes. In terms of point-to-mesh transformation, we utilize the Delaunay triangulation approach [107, 108] to convert the point cloud into meshes, which is demonstrated as a very effective method for surface reconstruction. It connects the points in point cloud scenes by forming triangles that satisfy the Delaunay criterion which guarantees no point lies inside the circumcircle of any triangle. This method generates a triangle mesh that approximates the surface of the point cloud [109]. In terms of mesh-to-image transformation, we leverage the rendering pipeline including the vertex transformation, projection, and rasterization, as well as shading. We directly use the rendering library OpenGL [110] to render images from the meshes. The process involves projecting the 3D vertices onto a 2D image plane based on camera parameters and applying shading and lighting calculations of 3D meshes to determine the specific color of each pixel. By this simple design, the world-to-camera extrinsic matrix $T_{e}$ containing both rotation and translation information between the 2D pixels and 3D points can be easily obtained.
2D to 3D Alignment. After multi-view rendering, the strict 2D-3D alignment can be easily established if the camera’s intrinsic $T_{i}$ is obtained from the standard calibration [111] and the extrinsic $T_{e}$ is obtained from the rendering. To be more specific, given the 3D point $\textbf{p}_{3D}\in\mathbb{R}^{3}$ as well as its 2D corresponding pixel coordinate $\textbf{p}_{2D}=(u,v)$, if we consider the pin-hole camera model, the transformation can be represented as $\hat{\textbf{p}}_{2D}=T_{i}~{}\cdot~{}T_{e}~{}\cdot\hat{\textbf{p}}_{3D}$. The $\hat{\textbf{p}}_{2D}$ and $\hat{\textbf{p}}_{3D}$ are represented within the homogeneous coordinates, and they are strictly paired. Therefore, we can strictly determine the correspondence between $\hat{\textbf{p}}_{2D}$ and $\hat{\textbf{p}}_{3D}$. Moreover, we can find an explicit association between each element of the textual feature $\textbf{F}_{T}$ and the 3D feature $\textbf{F}_{3D}$ while passing through the backbone network.
United 2D and 3D Proposal Generation. As shown in Fig. 6, according to our experiments, we found that it is difficult to directly find the object-level information merely based on the 2D proposals due to the great information loss while conducting rendering, the proposals provided by the 2D region proposal network (RPN) [112] can not effectively capture the 3D object information within the scene. On the other hand, merely relying on the 3D proposal provided by the 3D RPN [113] still can not guarantee accurate proposal generation for the fact that some objects are too adjacent in their geometry. To this end, we propose to leverage the union of 2D and 3D RPN to capture intact and all-inclusive proposals within the 3D scenes. Denote the region proposals as $R_{2D}$ and $R_{3D}$ respectively, the final holistic proposal generation $R_{H}$ is formulated as $R_{H}=R_{2D}\cup R_{3D}$. According to our experiments, this simple design can considerably boost the performance for the fact that the objects are more clustered and closely distributed within the indoor scenes. Taking the union of 2D and 3D object proposals into consideration also guarantees that the optimization merely considers the regions where the object really exists and prevents the models from taking the pure background into consideration during the optimizations.
Scene-level Sentence-3D Matching. We perform image-language alignment at both the object level and the scene level to pre-train the backbone network, which captures very accurate language-2D association for rendered images. The alignment at the object level allows the capturing of the fine-grained object-to-semantic concept matching, while the alignment at the scene level captures both holistic scene-level information and the dependent relations between the objects and holistic 3D scenes. For example, the bed exists in the bedroom while the sofa and TV lie in the living room with a clear owner-member relationship. To capture the object proposal information from the rendered images, we utilize the region proposal network (RPN) [112] to obtain the class-agnostic regions. For the text descriptor within the region, we use three most similar noun phase extract by it given by CLIP [9], which captures the scene level and the object level image patches with the most relevant texture description. Then we pass through both the original scene-level image as well as the obtained region-level image into the frozen CLIP visual encoder to obtain the scene-level visual feature $\textbf{I}_{\textbf{g}}$ as well as the object-level visual feature $\textbf{I}_{\textbf{l}}$, respectively. As shown in Figure 2, in order not to influence the well-learned feature representations provided by the CLIP encoder, we freeze the backbone and merely add a linear projection layer at the end of the network to facilitate the adaptive hierarchical feature-aligned pre-training.
Global and Regional Word-3D Matching. At the next step, we perform regional word-to-3D matching. We use the CLIP image encoder to generate a set of local feature embeddings $\textbf{F}^{l}_{3D}$ for the rendered proposal-level image patches, and the set of global embedding $\textbf{F}^{g}_{3D}$ for the whole scene. While the text encoder extracts the local textual embeddings $\textbf{F}^{l}_{T}\in\mathbb{R}^{H\times D}$ and the global textual embeddings $\textbf{F}^{g}_{T}$. The operation can be interpreted as a bi-directional operation, which means that for each region proposal image patch, we find the textual concept that matches best with the semantics of the region. And for each textual concept, we find the multiple regions that correspond with it. Next, we calculate the cosine similarity $\mathcal{G}_{sim}$ between the 3D visually encoded features and the textual features:
$\textbf{S}(\textbf{I},\textbf{T})=\mathcal{G}_{sim}(\textbf{F}^{l}_{3D},\textbf{F}^{l}_{T})+\mathcal{G}_{sim}(\textbf{F}^{g}_{3D},\textbf{F}^{g}_{T})$ (1)
[…]
Contrastive Optimizations. Note that when conducting contrastive learning, we regard textual features as anchors because textual descriptions are highly semantic and contain rich information, whereas the images contain too much low-level information and pixel-level details. Denote the $\textit{{F}}^{\;b,+}_{3D}$ and the $\textit{{F}}^{\;b,-}_{3D}$ as the positive and the negative features with respect to the anchor textual feature $\bm{F}_{T}^{a}$, respectively, the designed contrastive loss is formulated as follows:
[…]
$\mathcal{L}^{total}_{Ctr}=\mathcal{L}^{G}_{Ctr}+\mathcal{L}^{L}_{Ctr}$ (3)
3D Language Distillation. We utilize the $\mathcal{KL}$ divergence as the distillation loss to further distill the knowledge from 2D feature space to 3D. Compared with the mean square error loss such as the $\mathcal{L}_{1}$ or $\mathcal{L}_{2}$ losses, the $\mathcal{KL}$ divergence has improved regression capacity and ensured smoother gradient [115], which to some extent overcomes overfitting problems while distilling knowledge. The distillation loss $\mathcal{L}^{Dist}_{\mathcal{KL}}$ is formulated as:
[…]
According to our extensive experiments, our simple pre-training approach provides well-aligned vision-language-3D co-embedding. It largely facilitates vision-language-associated knowledge transfers from 2D to 3D, boosting both the label efficiency and the final recognition capacity of novel categories.
3.2 Region-Aware Fine-tuning
During the fine-tuning stage, our proposed framework consists of three subparts for the network optimization: 1. Unsupervised energy-based loss guided by boundary awareness and highly confident network predictions for unlabeled data, which is discussed in 3.2.1; 2. Unsupervised multi-stage region-level contrastive learning with highly confident predictions for unlabeled data, which is discussed in 3.2.2. 3. Supervised semantic contrastive learning for labeled data, which is discussed in 3.2.3. The three modules above are integrated jointly into the optimization function for network training to accomplish the final downstream detection or segmentation tasks with a very limited labeled data and a large amount of unlabeled data.
3.2.1 Unsupervised Region-level Boundary Awareness
Energy-function-based conditional random field segmentation is proposed in [29] and has been widely applied. However, it works in a fully supervised manner and does not explicitly consider the semantic boundary information, which is a great indication of semantic partitions in 3D scenes. In this Subsection, we develop a boundary-aware energy-based loss for unsupervised learning. As shown in Fig. 1, to obtain robust boundaries for unlabeled 3D points, we first perform 3D over-segmentation [34] and also extract boundary points using an off-the-shelf semantic boundary prediction network, which is both subsequently used as the conditions to define the boundary regions for 3D points. Then, we propose a region-level energy-based loss based on obtained boundary region labels.
Point Cloud Over-Segmentation. To obtain boundary regions, and facilitate subsequent region-level affinity computation and region-level contrastive learning, we first perform a region-level coarse clustering-based on point cloud over-segmentation. The previous method depended on the region growing [34] for over-segmentation, which relied heavily on the accurate normal estimation and was easily influenced by noises. In our work, we choose to use normal, curvature, and point feature histogram (PFH) [116] simultaneously to provide initial over-segmentation. The detailed procedures are given in the Appendix. And the over-segmentation results are shown in Fig. 1 and 5. Denote original point clouds as $P_{in}$. After over-segmentation, they are partitioned into $Q$ subregions $S=\{s_{1},s_{2},...,s_{q}\}$, where ${s_{i}}\cap{s_{k}}=\emptyset$ for any ${s_{i}}\neq{s_{k}}$ as shown in Fig. 1 and 5.
Boundary Points Extraction. As shown in Fig. 1, in addition to the over-segmentation result, we first extract the semantic boundary points to further identify boundary regions. The semantic boundary often indicates the distinguishment between various semantic classes. We extract semantic boundary points by JSENet [117], as shown in Fig. 3. Note that the JSENet is trained with diverse ratios of labeled data in weakly supervised manners. As for training, we first define semantic boundary points from the limited labeled scenes as ground truth. With the definition of the ground truth boundary points, we design the loss following JSENet except substituting the binary cross entropy loss $L_{bce}$ with the focal loss $L_{foc}$ [118] to tackle large class imbalances between the boundary points and non-boundary points. $L_{foc}$ can be formulated as follows:
$L_{foc}=-\frac{1}{N_{i}}\sum_{i=1}^{N_{i}}{(1-b^{gt}_{i})}^{\alpha}log(b_{i})+{(b^{gt}_{i})^{\alpha}}log(1-b_{i}),\vspace{-0.58 mm}$ (6)
where $b_{i}$ denotes the binary predicted boundary map and $b^{gt}_{i}$ denotes the ground truth boundary map. $N_{i}$ is the total number of input points for training. We select $\alpha$=2 based on the original design [118]. After its convergence, we apply the trained network to the remaining unlabeled scenes to obtain their boundary points. Examples of predicted boundary points of ScanNet [20] are shown in Fig. 5, which clearly reveal distinctions between diverse semantic classes.
Figure 4: Qualitative semantic segmentation results of the proposed WS3D++ on SemanticKITTI validation set the with the 5% labeling percentage, compared with the fully supervised state-of-the-art Cylinder3D [119], and BAAF-Net [89] with the semantics indicated by different colors. The red circles highlight the final performance difference between diverse approaches.
Boundary Labels. After extracting semantic boundary points, we utilize them as labels of discrimination between diverse semantic categories. As shown in Fig. 1, denote the $j_{th}$ adjacent regions of the center region $s_{i}$ as $s_{i,j}$. The adjacent region query is realized by fast Octree-based K-nearest neighbor search [120]. Then, we determine the two adjacent regions as boundary regions if both $s_{i}$ and $s_{i,j}$ contain boundary points. The label for boundary region $w_{i,j}$ is designed as:
$w_{i,j}=\begin{cases}1&\textit{if $s_{i}$, $s_{i,j}$ both contain boundary points;}\\ 0&\textit{otherwise.}\end{cases}$ (7)
The label $w_{i,j}$ denotes semantic boundaries of adjacent regions, which is then used to guide the optimization of the energy function for the downstream segmentation task.
Energy Loss Guided by Boundary Labels. As shown in Fig. 3, we first perform data augmentation (detailed in the Appendix) for the input point clouds $\textbf{P}_{in}$ to obtain two transformed point clouds $\textbf{P}_{in,1}$ and $\textbf{P}_{in,2}\in\mathbb{R}^{N\times C_{in}}$, where N is the numbers of points. $C_{in}$ and $C_{out}$ are the numbers of input and output feature channels, respectively. Utilizing the backbone network SparseConv [59], we can obtain point cloud predictions $\textbf{P}_{out,1}$ and $\textbf{P}_{out,2}$. Then applying region-level max pooling on the same subregions, we obtain the predicted classes $\textbf{P}_{1}$, and $\textbf{P}_{2}\in\mathbb{R}^{M\times C_{out}}$ of the specific subregions. $\textbf{P}=\{p(s_{1}),p(s_{2}),...p(s_{i}),...,p(s_{R})\}$, where $R$ is the total number of regions obtained by over-segmentation. Denote the prediction of the $j_{th}$ neighboring region of the center region $s_{i}$ as $p(s_{i,j})$. Taking the unary network prediction and pairwise affinity between the neighbouring region into account, inspired by conditional random field (CRF) in DeepLab [29], we formulate the optimization energy function $E_{sum}$ as follows for label-efficient learning:
$E_{sum}=\sum_{i}E_{i}(s_{i})+\sum_{i<j}^{adjacent}E_{i,j}(s_{i},s_{i,j}).$ (8)
The first unary network prediction item $E_{i}(s_{i})$ is the entropy regularization term. It encourages region-level prediction with high confidence, which also facilitates the contrastive learning introduced subsequently, which is formulated as:
$E_{i}(s_{i})=-log\,p(s_{i}).$ (9)
We propose the pairwise affinity term of $E_{sum}$ as:
$\begin{split}E_{i,j}(s_{i},s_{i,j})=H_{i,j}w_{i,j}[\epsilon-\|p(s_{i})-p(s_{i,j})\|]^{2}_{+}\\ +H_{i,j}(1-w_{i,j})\|p(s_{i})-p(s_{i,j})\|^{2}.\end{split}$ (10)
$H_{i,j}$ is designed as the confidence indicator. $H_{i,j}=1$, if the probabilities which produce $p(s_{i})$ and $p(s_{i,j})$ are both larger than a threshold $\gamma$. Otherwise, it equals 00. $\epsilon$ can be any value in the range of $(0,1)$. $p(s_{i})$ and $p(s_{i,j})$ are the semantic predictions for the $i_{th}$ center region and the $j_{th}$ neighbouring region, respectively. And $[x]_{+}$ is the maximum function $max(0,x)$. For adjacent boundary regions, we encourage their confident semantic or instance predictions to be different (i.e. larger $\|p(s_{i})-p(s_{i,j})\|$); while for the non-boundary adjacent regions, we force their semantic predictions to be the same (i.e. $\|p(s_{i})-p(s_{i,j})\|=0$). Different from the traditional energy function in DeepLab[29], which used handcrafted features to compare similarities, we propose to use the learned boundary region labels to guide the network’s confident max-pooled region-level predictions. Therefore, the proposed boundary-aware energy function better encourages semantic separations at boundaries. Furthermore, we only consider pairwise affinity between adjacent regions instead of all pixel pairs, which greatly reduces computation costs, and avoids noises induced by distant unrelated pairs in the meanwhile.
Figure 5: Qualitative instance segmentation results of proposed WS3D on ScanNet [20] with fully supervised state-of-the-art and the previous state-of-the-art GPC [121] in the 5% labeling ratio, compared with fully supervised arts. Different instances are indicated by different colors. And the intermediate over-segmentation results for obtaining regions and boundary predictions, with the boundary points indicated by the color blue.
3.2.2 Region-level Instance Discrimination
After applying the entropy regularization term $E_{i}(s_{i})$, we can obtain predictions at the region level with high confidence. Note that confident region-level predictions further improve the latent feature discrimination capacity of the network, which makes contrastive learning in the latent space feasible. Therefore, we further propose a multi-stage region-level contrastive learning for unlabeled data. Compared to previous work using only contrastive learning with low-level geometric registrations [27], our work unleashes potentials of contrastive learning with instance discrimination to enhance versatile feature learning in latent space.
The key to successful semantic/instance segmentation is maintaining discriminative feature representations at different stages of the backbone network [122]. Inspired by the pyramidal feature network [122, 123], we propose a simple but effective multistage contrastive learning approach for point clouds in an unsupervised setting. As shown in Fig. 3, given the augmented point clouds input $P_{in,1}$ and $P_{in,2}$, we feed them into the backbone encoder. We add five additional MLP heads with region-level max-pooling to obtain region-level segmentation predictions at the $m_{th}$ backbone stage, denoted as $\textbf{f}_{g,m}$ and $\textbf{f}_{h,m}$, respectively (five stages in our case, denote $M$ as the total network stages. i.e., $M$ = 5). After we apply the MLP heads to the extracted features at different stages, we can obtain the hierarchical feature embeddings of the fine-tuned downstream 3D scene understanding network. Unlike existing pixel-level [124], or point-level [124, 125] for point cloud contrastive learning, our proposed contrastive learning performs at the region level. Therefore, it will help to obtain more holistic regional information. The obtained regional information will be complementary to the point-level information. The region-level semantic contrastive loss is formulated as:
$L^{m}_{contrast}=-\frac{1}{\textbf{S}_{p}}\sum_{(a,b)\in S_{p}}log\frac{H_{i,j}exp(\textbf{f}_{g,m}^{a}\cdot\textbf{f}_{h,m}^{+}/\tau)}{\sum_{(\cdot,c)\in S_{p}}H_{i,j}exp(\textbf{f}_{g,m}^{a}\cdot\textbf{f}_{h,m}^{-}/\tau))},$ (11)
where $(a,b)\in\textbf{S}_{p}$ are latent confident predicted positive region pairs, and $(a,c)\in\textbf{S}_{p}$ are latent negative region pairs. As mentioned before, $H_{i,j}$ is designed to eliminate contrastive learning candidates with low confidence degrees. Reliable region-level contrastive learning is only applied to confident predictions by the network. After applying contrastive losses for instance discrimination, semantically similar features with high confidence levels in the latent space will become more adjacent, whilst semantically distinct ones will be separated, which will be shown in Fig. 13. Note that although a recent work GPC [121] has proposed methods to perform contrastive learning on the point clouds in a SSL manner, our work is different from their method in two aspects. Firstly, our contrastive learning is conducted at region level while GPC conducts contrastive learning at the point level. Secondly, GPC focuses on the selection of the positive and negative point-set samples to perform contrastive learning in a pseudo-label supervised manner on two different 3D scene samples, while we focus on the unsupervised contrastive learning which disentangles different feature representations in the latent space on the same augmented 3D scene sample, guided by confident network predictions, and it learns hirachical regional object information.
The final proposed multi-stage contrastive learning loss is formulated as the sum of losses at every network stage:
$L_{multi}=\sum_{m=1}^{M}L_{contrast}^{m},\vspace{-0.1mm}$ (12)
After applying the multi-stage contrastive loss $L_{multi}$, the output at each stage of the network will provide more distinctive representations to attain better performance. From our ablation experiments, the performances can be boosted by applying multi-stage contrastive loss. We apply the typical self-training-based approach, where the high-confidence predictions are retained as the pseudo label for the network training in the subsequent iteration. We adopt the typical cross-entropy loss for the self-training which is denoted as $L_{ST}$. Considering $E_{sum}$ (see Eq. 8) for the two augmented scenes $P_{in,1}$, $P_{in,2}$, and $L_{multi}$ (see Eq. 12), we formulate the overall loss $L_{unlabeled}$ for the WS3D training with unlabeled data: $L_{unlabeled}=L_{ST}+\alpha E_{sum,2}+\beta L_{multi}$.
3.2.3 Supervised Learning for labeled data
We also guide network optimizations by using supervision from the labeled data. As shown in Fig. 3, we use the cross-entropy loss $L_{ce}$ to guide the supervised learning on the labeled data in the supervised branch. The loss term for the WS3D training with the labeled data is $L_{labeled}$. For the task of semantic segmentation, the loss for the labeled data $L_{labeled}$ includes the cross-entropy loss. For the backbone network selection, we follow previous work with the sparse convolutional network design for semantic [59] and instance segmentation [90], respectively. Our detection network is designed based on widely adopted VoteNet [126, 99]. We also use Dice [127] loss to guarantee tighter aggregations of points within the same cluster, and strict geometric separations of points in diverse clusters [99].
3.2.4 The Overall Optimization Loss Function
Leveraging our proposed region-level energy-based loss and region-level contrastive learning, the network can make maximum use of the unlabeled data for better feature learning to boost performance in the fine-tuning stage. As shown in Fig. 3, for semantic segmentation and instance segmentation, we fine-tune the network in an end-to-end manner for both supervised and unsupervised branches to make full use of labeled and unlabeled data. The weight value of the unsupervised loss is denoted as $\lambda_{u}$. The overall optimization function $L_{total}$ is formulated as follows, and we set $\lambda_{u}=2$ empirically:
$L_{total}=L_{labeled}+\lambda_{u}L_{unlabeled}.$ (13)
Figure 6: The captured 2D and 3D region proposals. It is demonstrated that more object proposals can be captured by our proposed combined unitive 2D/3D proposal generation approach.
Figure 7: The final ScanNet [20] object detection results. Differences in predictions are highlighted in the rectangles.
Figure 8: The object detection comparisons on KITTI [128] validation set. It can be demonstrated that our proposed WS3D++ provides very accurate bounding box predictions qualitatively compared with the previous state-of-the-art merely using 2D bounding boxes, which demonstrates the effectiveness of using 3D regional as well as 3D object-level information in facilitating effective feature representation learning.
Figure 9: Segmentation result comparisons with CLIP prompts for the indoor ScanNet benchmark. It can be seen from our results that clear object-level information can be captured with the designed visual prompts, which demonstrates the effectiveness of our designed hierarchical feature-aligned representation learning approaches.
4 Experiments
4.1 Pre-training Experimental Settings
For the indoor scene understanding tasks, we pre-train the network on ScanNet [20]. And for the outdoor scene parsing tasks, we pre-train the network on NuScenes [35] dataset. For the dataset partition, we follow the official partition of ScanNet-V2 [20] using 1,201 scans as the pre-training dataset. The NuScenes [35] is an outdoor autonomous driving dataset that contains 7000 training scenes, the dataset provides the camera’s intrinsic and extrinsic parameters, thus we can obtain the 2D to 3D transformations and alignments very easily from designed rendering approaches. For the indoor and outdoor pre-training, we pre-train the network for 500 epochs and then we fine-tune the network on diverse downstream tasks. The hyper-parameter $k$ in Top-$k$ is set to 3. The initial learning rate is set to $5\times 10^{-4}$ and is multiplied with 0.2 every 50 epochs.
4.2 Finetuning Experimental Settings
Datasets. During the fine-tuning, to demonstrate the effectiveness for both data-efficient learning and open-world recognition of the proposed WS3D and WS3D++ under the limited scene reconstruction labeling scheme, we have tested it on various benchmarks, including S3DIS [21], ScanNet [129], and SemanticKITTI [18] for semantic segmentation, and ScanNet [129] for instance segmentation, respectively. Detailed information on each dataset and training details are put into the Appendix.
Training Set Partition. Following the typical setting in data-efficient learning in the limited reconstruction case [22] [121], we partition the training set of all tested datasets into labeled data and unlabeled data with various labeling points percentage, e.g., {1%, 5%, 10%, 15%, 20%, 25%, 30%, 40%, 100%}. For the limited reconstruction case, noted that to partition the labeled points into a specific labeling ratio, we probably need to split a maximum of one scene into two sub-scenes. One of the sub-scenes belongs to the labeled data and the other sub-scene belongs to the unlabeled data.
TABLE I: Comparison of semantic segmentation results with different labeling percentages on ScanNet validation set, S3DIS validation set (Area 5), and SemanticKITTI validation set (Sequence 08). ’Sup-only-GPC’ denotes GPC model trained with only labeled data. ’WS3D’ denotes model trained with our proposed methods. We have shown the performance increase in the last row for each dataset, compared to merely trained models with labeled data (left value) and to the SOTAs GPC [121] (right value).
Datasets Approaches Semantic Segmentation mIoU (%) on the Validation Set According to Supervision Level (%)
1% 5% 10% 15% 20% 25% 30% 40% 100%
ScanNet Sup-only-GPC 40.9 48.1 57.2 61.3 64.0 65.3 67.1 68.8 72.9
GPC [121] 46.6 54.8 60.5 63.3 66.7 67.5 68.9 71.3 74.0
Mix3D [130] 47.1 55.2 61.1 63.5 66.8 67.9 68.7 71.5 73.3
PointMixup [131] 44.5 55.2 69.6 63.7 66.3 66.9 68.8 70.2 73.8
RDPL [132] 47.6 56.3 61.1 63.8 66.8 67.8 69.7 71.8 74.6
Active-ST [133] 48.7 56.7 62.5 65.6 67.8 68.7 71.2 72.5 75.8
WS3D [4] 49.9 +9.0 56.2 +8.1 62.2 +5.0 65.8 +4.5 68.5 +4.5 69.4 +4.1 70.3 +3.2 73.4 +4.6 76.9 +4.0
WS3D++ 52.6 +11.7 59.1 +11.0 65.2 +8.0 66.9 +5.6 71.8 +6.6 71.1 +5.8 72.6 +5.5 75.3 +6.6 81.7 +8.8
S3DIS Sup-only-GPC 36.3 45.0 52.9 53.8 59.9 60.3 61.2 62.6 66.4
GPC [121] 38.2 53.0 57.7 60.2 63.5 63.9 64.9 65.0 68.8
Mix3D [130] 39.2 50.3 58.3 61.1 64.1 64.8 65.9 66.9 69.6
WS3D [4] 45.3 +9.0 54.6 +9.6 59.3 +6.4 62.3 +7.0 65.7 +5.8 66.5 +6.8 67.2 +6.0 69.5 +6.9 72.9 +6.5
WS3D++ 48.6 +12.3 57.7 +12.7 61.2 +8.3 66.9 +11.6 70.6 +10.7 71.1 +10.8 72.6 +11.4 75.3 +12.7 81.7 +15.3
SemanticKITTI Sup-only-GPC 28.6 34.8 43.9 47.9 53.8 55.1 55.4 57.4 65.0
GPC [121] 34.7 41.8 49.9 53.1 58.8 59.1 59.4 59.9 65.8
LESS [134] 37.1 42.5 50.5 53.9 59.5 59.6 60.5 63.5 66.9
Mix3D [130] 37.7 42.9 50.8 54.1 59.9 60.9 61.8 61.3 68.8
WS3D [4] 38.9 +10.3 43.7 +8.9 52.3 +8.4 55.5 +7.6 61.4 +7.6 61.8 +6.7 62.1 +6.7 63.2 +5.8 66.9 +1.9
WS3D++ 46.8 +18.2 58.6 +13.8 55.2 +11.3 62.8 +14.9 65.9 +12.1 67.9 +12.8 68.6 +13.2 70.9 +13.5 76.8 +11.8
Nuscene Sup-only-GPC 39.8 45.9 47.1 48.6 55.7 56.3 58.9 59.8 67.9
PRKT [135] 48.0 49.3 51.8 53.5 56.7 59.6 63.8 66.7 70.1
SLidR [134] 48.2 50.3 52.9 55.8 56.9 59.1 60.8 63.8 66.9
PointContrast [27] 48.3 50.2 55.1 58.1 60.2 61.1 62.1 64.9 69.2
CLIP2Scene [27] 56.3 56.8 58.7 61.2 63.6 64.1 64.5 64.8 65.1
LESS [134] 49.1 42.5 50.5 53.9 59.5 59.6 60.5 63.5 66.9
Mix3D [130] 37.7 42.9 50.8 54.1 59.3 60.2 61.8 61.3 68.8
WS3D [4] 49.1 +9.3 51.6 +5.7 52.3 +5.2 55.5 +6.9 61.4 +6.7 61.8 +5.5 62.1 +3.2 63.9 +4.1 71.6 +3.7
WS3D++ [4] 53.7 +13.9 56.7 +10.8 55.2 +11.3 62.8 +14.9 71.6 +15.9 67.9 +11.6 68.6 +9.7 71.9 +12.1 77.8 +9.9
Figure 10: The final 3D scene-level activations results based on the language query of WS3D++. Better zoom in for details.
[…]
4.3 Data-efficient 3D Semantic Segmentation
Overall Experimental Results. For the semantic segmentation, we have tested WS3D on versatile indoor and outdoor benchmarks, including ScanNet [20], S3DIS [21], and SemanticKITTI [18]. We have done extensive experiments with limited labeled data, e.g., only {1%, 5%, 10%, 15%, 20%, 25%, 30%, 40%, 100%} data in training set are available as labeled data. The qualitative results are shown in Fig. 4. In the meanwhile, the quantitative semantic segmentation performance is summarized in Table I. As mentioned, we have used the voxel-based method SparseConv [59] as the backbone. Our WSL model significantly surpasses the supervised-only model in GPC that is merely trained with labeled data, showing that our WSL can effectively make use of the unlabeled data to enhance the feature discrimination capacity of the model. Also, it can be observed the increment of performance is more obvious when the unlabeled data percentage is larger. For example, the performance increase on SemanticKITTI is 10.3% for the 1% labeling percentage, 5.8% for the 40% labeling percentage, and 1.9% for the 100% labeling percentage. This can be possibly explained by the fact that for more unlabeled data, our proposed WS3D can extract more meaningful semantic information from the unlabeled data based on our boundary-guided energy-based loss and confidence-guided region-level contrastive learning design. In addition, compared with current SOTA GPC, our proposed WS3D also achieves consistently better results in semantic segmentation performance, especially when faced with very limited label circumstances (e.g. 1% labeling points). In that case, WS3D outperforms GPC by 3.3%, 7.1%, and 4.2% for ScanNet, S3DIS, and SemanticKITTI, respectively. Fig. 4 shows that we can provide comparable performance compared with fully supervised SOTAs BAAF-Net [89] and Cylinder3D [119] on SemanticKITTI with 5% labels. As shown in Table I, the performance of our enhanced approach WS3D++ has remarkably increased performance compared with WS3D and previous SOTAs, demonstrating the effectiveness of our proposed vision-language knowledge-associated pre-training.
Figure 11: Segmentation result visualizations and comparisons with CLIP text encoder prompts for the outdoor KITTI benchmark. It can be demonstrated the foreground object awareness can be clearly captured comared with the previous SOTAs PLA [85]. Meanwhile, as shown in the last three columns, we can provide clear segmentation for the corresponding visual objects based on the textual prompts.
TABLE II: Comparison of experimental results on 20% and fully labeled case for the task of inductive and transductive learning for our proposed WS3D++, respectively. In transductive learning, the test set is also utilized for network training. We test the task of semantic segmentation on ScanNet, S3DIS, and SemanticKITTI with the evaluation metric of mIoU(%).
Datasets 20% label 100% label
Base Induct. Transduct. Base Induct. Transduct.
ScanNetv2-WS3D 64.0 68.5 71.4 72.9 76.9 77.6
ScanNetv2-WS3D++ 67.3 73.9 75.6 77.3 80.3 81.6
S3DIS Area5 Val. 59.9 65.7 66.6 66.4 72.9 73.5
S3DIS WS3D 61.3 67.6 68.6 68.2 74.8 75.8
S3DIS WS3D++ 62.8 68.9 69.9 69.7 75.5 76.9
Semantic KITTI Val. 53.8 61.4 64.5 65.0 66.9 68.2
Semantic KITTI Val. WS3D++ 54.7 62.9 66.2 67.1 68.2 69.7
Semantic KITTI Test. 55.7 62.5 63.6 65.4 68.1 71.3
Semantic KITTI Test. WS3D++ 56.6 63.7 65.3 67.1 69.7 72.8
TABLE III: Comparison of current State-of-the-art (SOTA) approaches in the limited reconstruction case for the 3D object detection tasks with different ratios of labeled data. The mean average precision (mAP) given by the mean $\pm$ standard deviation in three runs of diverse random splits are reported.
Datasets Models 5% 10% 20%
Induct. Transduct. Induct. Transduct. Induct. Transduct.
SUN RGB-D [137, 66] Baseline 29.9 $\pm$ 1.5 33.5 $\pm$ 0.8 34.4$\pm$ 1.1 40.7$\pm$ 0.9 41.1 $\pm$ 0.3 47.5 $\pm$ 0.5
SESS [138] 34.2 $\pm$ 2.0 38.1 $\pm$ 0.7 42.9 $\pm$ 0.8 45.3$\pm$ 0.9 47.9 $\pm$ 0.4 51.6 $\pm$ 0.3
3D-IOUMatch [139] 39.1 $\pm$ 1.9 46.3 $\pm$ 0.7 45.5 $\pm$ 1.5 53.5 $\pm$ 0.3 49.7 $\pm$ 0.5 54.3 $\pm$ 0.8
SPD [140] 38.5 $\pm$ 0.7 44.3 $\pm$ 0.8 46.0 $\pm$ 1.0 51.7 $\pm$ 0.9 49.6 $\pm$ 0.5 54.5 $\pm$ 0.9
WS3D (Ours) 46.9 $\pm$ 0.8 49.5 $\pm$ 0.7 47.8 $\pm$ 0.8 55.7 $\pm$ 0.8 55.6 $\pm$ 0.7 57.9 $\pm$ 1.2
WS3D-Open (Ours) 47.8 $\pm$ 0.8 50.3 $\pm$ 0.6 49.3 $\pm$ 0.7 58.2 $\pm$ 0.7 59.1 $\pm$ 0.8 60.3 $\pm$ 1.1
WS3D++ (Ours) 52.7 $\pm$ 0.7 54.5 $\pm$ 0.8 53.7 $\pm$ 0.9 63.8 $\pm$ 0.8 65.8 $\pm$ 0.7 65.6 $\pm$ 1.1
ScanNet-V2 [20] Baseline 27.9 $\pm$ 0.5 30.8 $\pm$ 1.5 31.1 $\pm$ 0.7 33.2 $\pm$ 0.5 41.6 $\pm$ 0.9 44.5 $\pm$ 1.2
SESS [138] 32.2 $\pm$ 0.8 36.8 $\pm$ 1.1 39.7 $\pm$ 0.9 44.5 $\pm$ 0.5 47.9 $\pm$ 0.4 49.2 $\pm$ 0.7
3D-IOUMatch [139] 40.0 $\pm$ 0.9 46.3 $\pm$ 0.7 47.2 $\pm$ 0.4 49.6 $\pm$ 0.6 52.8 $\pm$ 1.2 54.3 $\pm$ 0.8
SPD [140] 41.5 $\pm$ 0.5 44.3 $\pm$ 0.8 43.2 $\pm$ 1.2 46.2 $\pm$ 0.5 51.9 $\pm$ 0.5 55.1 $\pm$ 0.9
WS3D (Ours) 45.8 $\pm$ 0.5 49.2 $\pm$ 0.6 47.7 $\pm$ 0.9 53.9 $\pm$ 0.6 55.6 $\pm$ 0.7 58.2 $\pm$ 1.3
WS3D-Open (Ours) 47.2 $\pm$ 0.8 50.8 $\pm$ 0.6 51.1 $\pm$ 0.9 59.2 $\pm$ 0.8 59.0 $\pm$ 0.8 61.5 $\pm$ 1.6
WS3D++ (Ours) 55.3 $\pm$ 1.1 57.6 $\pm$ 1.2 57.9 $\pm$ 0.8 65.8 $\pm$ 1.2 64.6 $\pm$ 0.5 66.5 $\pm$ 1.1
TABLE IV: Comparisons of the open vocabulary learning performance. It can be demonstrated that our proposed approach provides very superior open-world recognition performance compared with the diverse SOTAs. WS3D-Open abandons our feature-aligned contrastive pre-training with the contrastive loss term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [9] feature encoder with the knowledge distillation loss term $\mathcal{L}^{Dist}_{\mathcal{KL}}$.
Datasets Models Few-shot settings
B15/N4 B12/N7 B10/N9
ScanNet [20] 3DGenZ [141] 20.6/56.0/12.6 19.8/35.5/13.3 12.0/63.6/6.6
3DTZSL [142] 10.5/36.7/6.1 3.8/36.6/2.0 7.8/55.5/4.2
LSeg3D [139] 0.0/64.4/0.0 0.9/55.7/0.1 1.8/68.4/0.9
PLA without caption [85] 39.7/68.3/28.0 24.5/70.0/14.8 25.7/75.6/15.5
PLA [85] 65.3/68.3/62.4 55.3/69.5/45.9 53.1/76.2/40.8
WS3D (Ours) 66.1/69.1/65.5 63.3/70.3/58.8 59.3/77.5/51.6
WS3D-Open (Ours) 68.3/70.6/67.1 64.2/71.4/59.9 59.7/77.9/51.6
WS3D++ (Ours) 72.3/69.1/73.3 73.3/70.0/63.6 65.2 / 78.9 / 58.8
Fully-supervised 74.5/68.4/79.1 73.6/72.0/72.8 69.9/75.8/64.9
Datasets Models Few-shot settings
B12/N3 B10/N5 B6/N9
NuScenes 3DGenZ [141] 01.6/53.3/00.8 01.9/44.6/01.0 01.1/52.6/00.5
3DTZSL [142] 01.2/21.0/0.6 06.4/17.1/03.9 2.61/18.52/03.15
LSeg-3D [85] 0.6/74.4/0.3 0.0/72.5/0.0 2.66/69.72/0.21
PLA without caption [85] 25.5/75.8/15.4 10.7/76.0/05.7 8.95/65.83/6.32
PLA [85] 47.7/73.4/35.4 24.3/73.1/14.5 15.63/60.32/12.38
WS3D (Ours) 51.3/76.6/30.9 41.5/72.5/20.8 35.2/40.9/29.8
WS3D-Open (Ours) 55.8/77.5/34.9 47.2/69.3/50.9 39.2/45.3/41.6
WS3D++ (Ours) 58.6/79.1/38.6 49.7/67.3/53.9 40.6/46.6/51.9
Fully-supervised 73.7/76.6/71.1 74.8/76.8/72.8 74.6/75.9/72.3
Figure 12: The WS3D++ corresponding detection and segmentation projections on six rendered 2D views through proposed multi-view rendering. It can be demonstrated that our proposed rendering has established an explicit modality association between the final 2D views and 3D views. Also, by combining 2D and 3D region proposals, more complete and apparent object-level information can be clearly captured both from 2D views and 3D views. Best zoom-in for viewing.
Figure 13: t-SNE visualization in semantic segmentation of the proposed WS3D under the 5% labeling percentage on S3DIS validation set. The diverse feature embeddings are indicated by different colors and are normalized into $[-1,1]$ for better visualization. It has been validated that more discriminative features can be acquired for diverse semantic classes with our proposed unsupervised region-aware fine-tuning strategy (demonstrated by WS3D) and our proposed hierarchical vision-language knowledge associated and distilled pre-training (demonstrated by WS3D++).
TABLE V: Comparison of the performance of instance segmentation, under various levels of supervision on ScanNet validation set. ’Sup-only-GPC’ denotes the model trained with only labeled data. ’WS3D’ denotes the model trained with our proposed methods. In the last row, we have shown the performance increase of WS3D. WS3D-Open abandons our feature-aligned contrastive pre-training with the contrastive loss term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [9] feature encoder with the knowledge distillation loss term $\mathcal{L}^{Dist}_{\mathcal{KL}}$.
Tested Dataset Approaches Ins. Seg. Results with the metric of AP@50%
1% 5% 10% 15% 20% 30% 35% 40% 100%
ScanNet Sup-only-GPC [121] 10.8 33.6 42.8 45.3 48.2 49.0 49.5 50.2 56.8
Mix3D [130] 12.7 34.7 43.1 45.7 48.7 49.6 50.2 51.3 57.6
GPC [121] 16.9 38.6 44.9 47.2 48.6 49.7 51.2 52.0 57.7
SPIB_Ins [143] 17.1 38.9 45.3 47.9 48.9 50.5 51.6 52.7 57.6
WS3D (Ours) [4] 30.8 45.6 53.5 54.7 53.2 51.9 52.5 53.0 58.7
WS3D-Open (Ours) 31.2 50.6 55.8 57.9 56.8 58.7 58.9 59.5 65.1
WS3D++ (Ours) 39.2 54.8 57.3 58 +8.6 +9.7 +9.4 +9.3 +8.3
4.4 Data-Efficient 3D Instance Segmentation
As our method can be integrated seamlessly into various network backbones and applied to different highly-level understanding tasks, we have also integrated our method with Point-Group [90] for the instance segmentation on ScanNet with results shown in Table V. Notice that the performance increase is 21.7% when merely 1% data is labeled compared with the sup-only case. It further demonstrates that our proposed approaches for the unsupervised branch have effectively exploited the unlabeled data to improve the feature learning capacity of the model. As shown in Fig. 5, our proposed WS3D and WS3D++ both provide explicit boundary guidance for separating diverse kinds of semantic classes, and the instance segmentation performance with very limited labeling percentage is comparable to those fully supervised counterparts.
4.5 Data-Efficient 3D Object Detection
For the data-efficient 3D object detection, following our previous work [99], we evaluate current approaches extensively on SUN RGB-D [137] and ScanNet [20] benchmarks for 3D object detection tasks with the strong 3D object detection backbone VoteNet [126]. It can be demonstrated the open-vocabulary designs can to some extent boost the 3D object detection performance, which demonstrate the generalization capacity of our fundation model. As shown in Table III, we have also tested the ablated approach WS3D-Open which abandons our feature-aligned pre-training term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [9] feature encoder for the knowledge distillation loss term with $\mathcal{L}^{Dist}_{\mathcal{KL}}$. It can be demonstrated that the performance degradation can obviously be observed when comparing WS3D-Open to WS3D, which demonstrate the effectiveness of our proposed hierarchical feature aligned pre-training in improve data efficiency.
TABLE VI: WS3D and WS3D++ ablation studies on ScanNet (Left Value), S3DIS (Middle Value) and Nuscene (Right Value) validation set, for semantic segmentation (Metric: mIoU%) and on ScanNet validation set for the instance segmentation (Metric: AP@50%), both tested with the 5% labeled case.
Cases Base $w_{i,j}$ $H_{i,j}$ in EF $H_{i,j}$ in UCSL UCSL MS-UCSL SCE WS3D mIoUWS3D++ mIoUWS3D AP@50% WS3D++ AP@50% WS3D mAPWS3D++ mAP%
No. 1 56.2 / 54.6 / 53.9 62.8 / 58.1 / 56.3 45.6 54.8 45.8 58.9
No. 2 51.0 / 49.3 / 47.8 57.1 / 52.5 / 50.9 39.9 48.5 40.6 53.6
No. 3 49.9 / 47.2 / 46.6 53.1 / 50.9 / 47.9 37.0 44.5 38.1 51.7
No. 4 51.6 / 52.1 / 47.9 56.1 / 54.8 / 47.7 40.1 46.7 41.6 51.5
No. 5 51.1 / 51.4 / 49.6 57.1 / 53.8 / 49.8 40.7 47.6 40.8 53.6
No. 6 52.5 / 50.9 / 50.3 56.7 / 53.9 / 49.5 42.2 48.7 39.8 53.8
No. 7 49.3 / 48.0/ 47.6 55.6 / 53.5/ 50.8 38.1 46.8 39.8 54.3
No. 8 54.3 / 52.8 / 50.8 60.6 / 55.3 / 52.7 42.9 48.1 41.2 52.7
No. 9 48.1 / 45.3 / 45.7 50.8 / 47.6 / 50.6 34.8 39.8 33.9 45.3
TABLE VII: Ablation studies of the designed models of our proposed WS3D++. It is demonstrated that the proposed proposal generation approach and local and global contrastive approach can significantly boost final recognition performance.
Cases 2D Proposal 3D Proposal $\mathcal{L}^{G}_{Ctr}$ $\mathcal{L}^{L}_{Ctr}$ mIoUAP@50% ScanNet B10/N9 Nuscene B6/N9
No. 1 62.7 / 58.1 / 56.3 58.5 / 55.6 / 53.7 65.2 / 78.9 / 58.8 69.7 / 71.7 / 66.8
No. 2 55.2 / 51.9 / 50.6 51.3 / 48.8 / 47.9 56.9 / 70.7 / 53.8 66.1 / 66.6 / 64.5
No. 3 53.6 / 50.6 / 49.6 52.6 / 48.5 / 47.6 52.6 / 48.5 / 47.6 53.6 / 48.5 / 47.6
No. 5 53.2 / 51.7 / 49.3 50.2 / 46.1 / 45.8 50.8 / 45.8 / 45.5 50.8 / 46.7 / 45.8
No. 6 51.2 / 49.2 / 49.1 49.2 / 45.8 / 45.2 48.8 / 44.7 / 44.2 48.2 / 43.8 / 43.2
4.6 Ablation Study
Ablations: The results for WS3D are summarized in Table VI. We have ablated network modules in all combinations of settings as follows. Take the ScanNet instance segmentation at AP@50% as examples: Case 1: The full WS3D. Case 2: Removing the boundary prediction network, and not using the guidance of $w_{i,j}$. The framework still consists of the supervised branch, unsupervised guidance of the energy function based on the predicted confident pseudo label, and contrastive learning. This setting leads to a significant drop of 5.7% on AP. Case 3: Removing the pairwise term in the energy-based optimization function $E_{sum}$, the AP drops largely by 8.6%. Case 4: Removing $H_{i,j}$ in the energy function, the performance drops by 5.5%. Case 5: Removing $H_{i,j}$ in the unsupervised contrastive learning, the performance drops by 4.9%. Case 6: Conducting contrastive learning only with the region-level feature $\textbf{f}_{g,5}$ and $\textbf{f}_{h,5}$ at the fifth network stage, rather than at multiple stages. The performance drops by 3.4%. Case 7: Removing the unsupervised region-level contrastive learning branch, the performance drops largely by 7.5%. Case 8: Removing the supervised learning branch with the cross-entropy loss, the performance drops by 2.7%.
The results for WS3D++ are summarized in Table VII. Case 9: Only using the supervised branch, the ins. seg. performance drops significantly by 10.8%. Case 10: As demonstrated in Table VII, dropping the 2D or 3D proposals will both result in the performance drops both for the semantic segmentation as well as instance segmentation. Case 11: As shown in Table VII, either removing the global contrastive loss or the local contrastive loss results in less effective contrastive representations and affects the final downstream scene understanding performance.
Analyses: From the above ablations, some important findings are summarized: For WS3D, firstly, not using our designed modules results in a significant performance drop (Cases No. 3, No. 7, and No. 9), which demonstrates the effectiveness of the proposed unsupervised branch and learning strategies to leverage the unlabeled data. Secondly, our proposed learning strategies with boundary label $w_{i,j}$ (Case No. 2), energy function design (No. 3), high-confidence prediction based energy function design (No. 4), high-confidence based region-level contrastive learning strategy (No. 5), multi-stage contrastive learning network design (No. 6), all have a boost on the overall semantic/instance segmentation performance. The results demonstrate that the proposed energy loss is significant for semantic/instance seg. performances, because semantic boundary labels are crucial for identifying diverse objects. Thirdly, removing the supervision (Case No. 8), our method still maintains performance with a slight drop of performance by 2.7%. It further validates the robustness and feature learning capacity of our approach. Last but not least, as demonstrated in Table VII, for our proposed WS3D++, either dropping the 2D/3D proposals, or dropping the global/local contrastive loss result in considerable performance drops. It validates the effectiveness of our designs in enhancing discriminative 3D representations.
4.7 Qualitative and Quantitative Results of the Open-world Recognition Approaches
In this Subsection, we further evaluate the performance of the open-world recognition capacity of our proposed approach. The results of open-world recognition are shown in Table IV. We have compared our work with the previous approach PLA [85] in establishing the sufficient point-language associations for the open-world robot learning. The results demonstrate that our proposed approach has superior performance in open-world recognition. We directly use the settings in the PLA [85] and split the categories on ScanNet [20] and Nuscene [35] into base and novel categories. It can also be validated that WS3D-Open, which abandons our feature-aligned pre-training and directly use the CLIP [9] feature encoder, provides slightly inferior performance compared with WS3D++, validating the effectiveness of our language-3D matching strategy designs. WS3D++ exhibits superior performance in terms of the open-vocabulary few-shot learning for diverse partitioning of original and novel classes. The open-world recognition results are shown in Figure 9 and Figure 11. It can be demonstrated that better foreground object awareness can be effectively capture by our proposed WS3D++ compared with PLA [85], with superior segmentation performance guided by the textual prompts. The superior open-world recognition performance can be achieved while conducting open-world learning in diverse spliting of based and novel classes, including B15/N4, B12/N7, B10/N9 for the ScanNet [20] as well as B12/N3 and B10/N5 for the NuScenes [35]. It demonstrates robustness of our proposed approach. Also, as demonstrated in Figure 10, the WS3D++ language driven-3D scene segmentation results are very precise as shown qualitatively, which is corresponding to the object queried by the language, and it demonstrates that the inference can be done based on the object, material, properties, affordance, room type, etc. It demonstrates that our proposed WS3D++ can enable the scene-level object recognition based on the semantic language queries. As further shown in Figure 12, through our effective rendering techniques, which establish the explicit 2D-3D association, the aligned representation of 2D-3D-language co-embeddings can be learned and the objectness can also be enhanced through finding the similarity among diverse views through contrastive learning approaches. Also, by combining 2D and 3D region proposals, more complete and apparent object-level information can be clearly captured both from 2D views and 3D views.
4.8 Instance Discrimination Capacity
We show t-SNE visualizations of the learned latent feature representations for various semantic classes in Fig. 13. The case study task is the semantic segmentation on the S3DIS dataset with a supervision level of 5%. It is demonstrated that more distinctive and better separated point-wise feature embeddings are provided by our proposed unsupervised region-level contrastive learning, which can be attributed to its strong instance discrimination capacity. And more separated feature space can be provided with our proposed WS3D++ compared to WS3D and GPC. This strong instance discrimination capacity can be explained by more discriminative feature representations guided by 3D vision-language aligned representations, and is thus more beneficial to high-level semantic and instance segmentation performances both in terms of data efficiency and open-world recognition capacity. Also, our proposed hierarichical feature alignment also provides more separated feature space, which means that the feature alignment successfully enhances the final instance discrimination capacity.
5 Conclusion
In this paper, we proposed a general WS3D++ framework for data-efficient 3D scene parsing. It involves both the pre-training and the fine-tuning stages. In the pre-training stage, we propose the hierarchical feature alignment strategy to acquire accurate regional 3D-linguistic pairs, thus the performance can be enhanced to a large extent. At the same time, we propose an unsupervised boundary-aware energy-based loss and a novel region-level multi-stage semantic contrastive learning strategy, which are complementary to each other to make the network learn more meaningful and discriminative features from the unlabeled data. The effectiveness of our approach is verified across three diverse large-scale 3D scene understanding benchmarks under various experiment circumstances. Our approach can maximally exploit the unlabeled data to enhance the performance both for 3D point clouds semantic segmentation and instance segmentation, under various labeling percentages in the limited reconstruction case. Our proposed label-efficient learning framework, termed WS3D++, provides conprehensive baselines for future 3D scene parsing methods when the label is inaccessible or limited.
Robot Arm Grasping Example
We deploy our approach for the task of open-world perception in robot grasping. We use our proposed approach for segmentation and use the ROS 2 Gazebo-based framework to implement the other components of the system, such as kinematics/dynamics modeling, motion planning, low-level control, point cloud-based pose estimation, etc. Our proposed approach has robust performance and decent accuracy in grasping, which demonstrates the potential of our proposed WS3D++ in industrial manipulation applications in grasping and dropping novel objects beyond the training set. The video demos are attached.
Supplementary Material
In this supplementary material, additional experimental results and details that are not included in the main paper due to space limits are provided as follows:
  • Details of our experimental datasets (see Section A).
  • Additional qualitative experimental results (see Section B).
  • The PFH-based point cloud over-segmentation method to provide the initial over-segmentation results for our proposed WS3D (see Section C).
  • Details of our data augmentation approach (see Section D).
  • Training details of the boundary prediction network (see Section E).
  • The low density issues for the outdoor point clouds captured by LiDAR sensors (see Section F).
A Experimental Datasets Details
To demonstrate the effectiveness of our proposed WS3D and WS3D++, we tested it on both indoor and outdoor 3D scene segmentation benchmarks, including S3DIS[21], SemanticKITTI[18], and ScanNet[20] for semantic segmentation, and ScanNet[20] for instance segmentation. The detailed information of each dataset is listed as follows:
Figure 14: Qualitative semantic segmentation results on SemanticKITTI with 5% labels, compared with the fully-supervised arts Cylinder3D [119], and BAAF-Net [89] with semantics indicated by diverse colors. The red circles highlight the performance difference between diverse methods.
ScanNet-V2 is a standard large-scale and widely-acknowledged 3D indoor scene understanding benchmark consisting of more than 1,600 3D scene scans. It contains densely reconstructed point clouds from RGB-D images captured by the depth camera. For the dataset partition in fully supervised learning, we follow the official partition of ScanNet-V2[20] using 1,201 scans as the training set, 312 scans as the validation set, and 100 scans as the test set.
SemanticKITTI[18] is a large-scale point cloud understanding benchmark for self-driving applications. The point cloud of individual scenes is obtained densely by Velodyne VLP-64 LiDAR. The dataset covers long-range road scenes, thanks to the vehicle driving in complex road scenarios with mounted LiDAR to capture the point clouds. More than 43,550 labeled LiDAR scans in total are split into 21 sequences. Each LiDAR scan contains approximately $10^{5}$ points. We follow the official setting[18] that uses the sequences 00 to 07, and 09 to 10 for training, the sequence 08 for validation, and the sequence 11 to 21 for testing.
S3DIS is a commonly used indoor 3D scene understanding benchmark that contains various indoor scenes. It includes six indoor areas which are made up of 271 rooms. Each room contains a magnitude of $10^{6}$ points. The typical room size is $20\times 15\times 5$ meters. In our experiments, we follow GPC[121] using area 5 as the validation set and the other areas as the training set.
NuScene is a commonly used autonomous driving dataset containing more than 1000 scenes. Every scene is elaborately annotated with the 3D bounding boxes for 23 categories and 8 attributes. For the base/novel category partition, we overlook the ’otherflat’ class and randomly divide the rest 15 classes into B12/N3 and B10/N5 as in PLA [85].
Figure 15: Results of our proposed PFH-based over-segmentation in Algorithm 1. Diverse over-segmented regions are indicated by diverse colors.
B Additional Experimental Results
We provide the additional experimental qualitative results on ScanNet and SemanticKITTI semantic segmentation and ScanNet instance segmentation.
In Fig. 14, the results on SemanticKITTI[18] are shown for the 5% labels case compared with the fully-supervised SOTAs Cylinder3D[119], and BAAF-Net[89]. From the qualitative results, our method can sometimes offer better segmentation performance than fully-supervised SOTAs for objects such as roads, trees, etc., which demonstrates the effectiveness of our proposed WS3D for outdoor 3D semantic segmentation.
Comparisons with SOTAs in the fully supervised mode. To demonstrate the feature learning capacity of our proposed WS3D, we have also experimented with the 100% labeling percentage for a fair comparison with fully supervised SOTAs. The results are shown in Table VIII and the last column of Table I. We have fed the whole training set into the supervised branch and unsupervised branch in Fig. 3 simultaneously. Therefore, our proposed energy-based loss and region-level contrastive learning strategy operate as additional optimization guidance for network training. Table VIII demonstrates that we can realize at least comparable or even better results, compared with the fully supervised SOTAs.
Transductive learning. Similar to the experimental setting of [121], we have also conducted experiments evaluating the performance of WS3D in transductive learning. Different from the inductive learning we tested above that requires the trained model to be generalized to an unseen test set, transductive learning can exploit the testing set when training. Compared with inductive learning, we add the test set as part of the unlabeled data in transductive learning. As is demonstrated in Table III, the sem. seg. mIOU becomes higher if the network is learned in a transductive way, both for the fully labeled case with 100% labels and the weakly labeled case with 20% labels. It demonstrates that our proposed WSL approaches, including the energy-based boundary-aware loss and the region-level contrastive learning, can leverage the unlabeled data for feature learning effectively in an implicit way to enhance the final segmentation performance.
Figure 16: In the limited reconstruction case, our proposed approaches WS3D and WS3D++ work well for the low-density SemanticKITTI over-segmentation and boundary prediction (5% label percentage). It can also be demonstrated that the explicit boundary information can be clearly captured, and our proposed approach is robust to the density variation of indoor dense and outdoor sparse point clouds.
C Point Clouds Over-segmentation
C.1 PFH-based Point Clouds Over-segmentation
As mentioned in Subsection 3.1 in the main paper, the previous over-segmentation methods depended on the region growing[34], which relied heavily on the accurate normal estimation and were easily influenced by noises. In our work, we use the normal, curvature, and point feature histogram[116] simultaneously to provide the initial over-segmentation results for obtaining different regions. The point feature histogram (PFH)[116] can be regarded as a robust representation of local geometric features, which is robust to 6D pose transformation and deals well with diverse sampling densities, and neighboring noises. We use it as a more robust geometric feature to obtain the over-segmentation result.
To be more specific, we denote the average normal vector of a point as $\textbf{n}=(n_{x},n_{y},n_{z})$, the PFH feature as $\textbf{f}_{pfh}\in\mathbb{R}^{125\times 1}$. The cosine of the angle between the average normal vectors of two adjacent points $P_{i}$ and $P_{j}$ is denoted as $A_{n}$, and the cosine of the angle between two PFH vectors of two adjacent points in the 125-dimensional space is denoted as $A_{pfh}$. The Affinity $A(P_{i},P_{j})$ of two adjacent points $P_{i}$ and $P_{j}$ is given as:
$A(P_{i},P_{j})=\sqrt{A_{n}^{2}+A_{pfh}^{2}}.$ (14)
According to the Point Cloud Library (PCL)[34], in the process of region growing, points with an affinity larger than a threshold $T_{thres}$ are merged iteratively. We set $T_{thres}$ as 1 following the typical setting[34]. As shown in Algorithm 1, we summarize the detailed procedures for the PFH-based over-segmentation as the following steps:
Input: The input raw point set $\textbf{P}_{i}=\{p_{i}\},i=1,2,...,N_{i}$. $p_{i}=(x_{i},y_{i},z_{i})$.
Output: The output pseudo label matrix $\textbf{{L}}_{region}$ for different regions.
Initialize $P_{seed}=p_{\{seed,i\}},i=1,2,...,N^{seed}$;
Initialize the pseudo label matrix $\textbf{{L}}_{region}$ as a zero matrix. while not converged do
       Select $K$ nearest neighbour points $p_{j}$ around the seed points $p_{seed}$ for comparisons based on fast Octree-based $K$ Nearest Neighbour search;
       for the seed points $p_{j}$ selected do
             if Condition 1 then
                  Assign $p_{j}$ the same overseg-label as $p_{seed}$;
                   if Condition 2 then
                        Regard the point $p_{j}$ as new seed points;
            else
                  Assign $p_{j}$ with a new overseg-label. Regard $p_{j}$ as new seed points;
            $j\leftarrow j+1$;
             Update the pseudo label matrix $\textit{{L}}_{region}$.
return The class label matrix $\textit{{L}}_{region}$ of $\textbf{P}_{i}$ with different clusters.
Algorithm 1 The Proposed PFH-based Region Growing Over-segmentation to Obtain Regions
  1. 1.
    Select the points that have top 2‰ minimum curvatures[34] as the initial seed points $p_{seed}$. Utilizing fast $K$ Nearest Neighbour search[120], we obtain the $j_{th}$ neighbouring point $p_{j}$ of the initial seed point $p_{seed}$, and calculate the Affinity $A(p_{seed},p_{j})$ between $p_{seed}$ and $p_{j}$.
  2. 2.
    If the Affinity $A(P_{seed},P_{j})$ is larger than $T_{thres}$ (Condition 1), we assign the $j_{th}$ neighbouring point $p_{j}$ with the same label as $p_{seed}$.
  3. 3.
    If Condition 1 is satisfied, denote the curvature of $p_{j}$ and $p_{seed}$ as $r_{j}$ and $r_{seed}$, and denote the difference of $r_{j}$ and $r_{seed}$ as $\Delta r=\|r_{j}-r_{seed}\|$. If $\Delta r\leq\sigma$ (Condition 2), we set $p_{j}$ as the seed point. That means we set the points $p_{j}$ as seed points, only if the difference in curvature is small enough. Otherwise, we only assign $p_{j}$ the same label as $p_{seed}$.
  4. 4.
    If Condition 1 is not satisfied, we assign the neighbouring point $p_{j}$ with a different label from $p_{seed}$, and also assign $p_{j}$ as the seed point.
Finally, points are assigned to initial over-segmented regions with labels. The simple algorithm is summarized in Algorithm 1. The loop will terminate if any one of the following convergence conditions is satisfied, which are proposed:
  1. 1.
    All points have been assigned with labels;
  2. 2.
    There are no seed points that can be added.
After the procedures above, we can obtain the label $\textit{{L}}_{region}$, which indicates the separated regions obtained by over-segmentation. Some randomly selected over-segmentation results are shown in Fig. 15. The PFH-based over-segmentation can automatically divide the whole point cloud scene into geometrically separated regions as shown in Fig. 15. Different colors indicate different regions.
TABLE VIII: Comparison of SOTAs methods in the semantic segmentation performance on ScanNet validation set, S3DIS validation set (Area 5), and on SemanticKITTI validation set and test set. All results are based on the 100% label ratio. Top-two results are highlighted.
Approaches Venue ScanNet S3DIS SemanticKITTI
Val. Val. (Area 5) Val. Test
Minkow-Network [37] CVPR19 72.2 65.4 61.1 63.1
PointASNL [49] CVPR20 66.4 62.6 - 58.8
KPConv [144] ICCV19 69.2 67.1 - 46.8
SPV-NAS [145] ECCV20 - - 64.7 66.4
Fusion-Net ECCV20 - 67.2 63.7 61.3
MV-Fusion [44] ECCV20 76.4 65.4 - -
Cylinder3D [119] CVPR21 - - 65.9 67.8
BAAF-Net [89] CVPR21 - 72.1 61.2 59.9
Sup-only-GPC [121] ICCV21 70.9 66.4 65.0 65.4
GPC [121] ICCV21 74.0 68.8 65.8 67.7
WS3D ECCV22 76.9 72.9 66.9 69.0
WS3D++ - 77.6 73.5 67.7 69.8
D Data Augmentation Details
D.1 Data Augmentation Approaches
We perform three kinds of data augmentations to enhance the unlabeled data and to perform region-level contrastive learning, including: 1. Geometric rotations and mirroring of the point cloud scene. 2. Adding minor disturbance to the point cloud scene. 3. Random scaling of the point cloud scene.
[…]
Adding minor disturbance. We add the noise for the normalized point clouds coordinates, which can be represented as $P_{x}^{d}=P_{x}^{m}+T_{d}$. $T_{d}$ is the Gaussian noise with the mean of 0 and variance of 0.01. We set $T_{d}$ in the range of $[-0.1,0.1]$.
Random scaling. We do random scaling to the point cloud scene. Random scaling of ($0.9-1.1$) is done to the point clouds coordinates. Denote the scaling parameter as $\eta$, the final augmented point clouds scene $P_{x}^{{}^{\prime}}$ is given as $P_{x}^{{}^{\prime}}=\eta P_{x}^{d}$.
E Training Details of the Boundary Prediction Network
[…]
F Low-Density Issues of Outdoor LiDAR points
While using the boundary prediction network to extract the semantic boundary for large-scale 3D scenes, the density issue should be taken into consideration because the density variation may influence the quality of the obtained pseudo labels. Randomly selected results of SKITTI LiDAR over-segmentation and boundary predictions results are shown in Fig. 16. The density influences over-segmentation performance especially for outdoor low-density point clouds because the 3D point cloud normals, curvatures, and PFH features are all directly related to density. For outdoor sparse SemanticKITTI, we have a very simple but effective trick to tackle low-density outdoor LiDAR points: discarding all regions containing less than $N_{thres}$ points, i.e. not using these regions in all our network modules. ($N_{thres}=100$ was set empirically for SKITTI). As shown in Fig. 16, isolated regions with fewer points than $N_{thres}$ are successfully discarded for reliable and robust region-level predictions. As shown in Fig. 16, compared to Karpathy’s approach, our PFH-based over-segmentation better distinguishes similar semantic classes such as road and sidewalk, revealing that our adapted PFH can have a better local feature description capacity than other classical features, both for our proposed WS3D and WS3D++. The results of boundary predictions are not much or greatly influenced by density as shown in Fig. 16. Our boundary prediction network learns high-level semantic boundaries, which are not explicitly related to the low-level point density. In summary, our proposed WS3D and WS3D++ still work well for the low-density cases.

### v2

@@
-Generalized Label-Efficient 3D Scene Parsing via Hierarchical Feature Aligned Pre-Training and Region-Aware Fine-tuning
+Generalized Robot 3D Vision-Language Model with Fast Rendering and Pre-Training Vision-Language Alignment
 Abstract
 Deep neural network models have achieved remarkable progress in 3D scene understanding while trained in the closed-set setting and with full labels.
-However, the major bottleneck for the current 3D recognition approach is that these models do not have the capacity to recognize any unseen novel classes beyond the training categories in diverse kinds of real-world applications.
+However, the major bottleneck for the current 3D recognition approach is that these models do not have the capacity to recognize any unseen novel classes beyond the training categories in diverse real-world applications.
 In the meantime, current state-of-the-art 3D scene understanding approaches primarily require a large number of high-quality labels to train neural networks, which merely perform well in a fully supervised manner.
 Therefore, we are in urgent need of a framework that can simultaneously be applicable to both 3D point cloud segmentation and detection, particularly in the circumstances where the labels are rather scarce.
-This work presents a generalized and simple framework for dealing with 3D scene understanding when the labeled scenes are quite limited.
+This work presents a generalized and straightforward framework for dealing with 3D scene understanding when the labeled scenes are quite limited.
 To extract knowledge for novel categories from the pre-trained vision-language models, we propose a hierarchical feature-aligned pre-training and knowledge distillation strategy to extract and distill meaningful information from large-scale vision-language models, which helps benefit the open-vocabulary scene understanding tasks.
@@
 Extensive experiments with both indoor and outdoor scenes demonstrated the effectiveness of our approach in both data-efficient learning and open-world few-shot learning.
-All codes, models, and data are to made publicly available at: https://github.com/KangchengLiu.
 The code is at: WS3D++ Code link.
@@
 1 Introduction
-The 3D scene parsing problem, which typically encompasses several important downstream tasks: point cloud semantic segmentation, instance segmentation, and object detection, becomes increasingly important with the wide deployment of 3D sensors, such as LiDAR and RGB-D cameras [1, 2, 3, 4, 5, 6, 7, 8].
+Figure 1: The final overall illustrative diagram of our proposed WS3D++.
+We integrate language-3D feature associated pre-training and data-efficient fine-tuning as a general scene parsing vision-language model to achieve effective data-efficient as well as open-vocabulary 3D scene understanding for 3D scenes.
+The 3D scene parsing problem, which typically encompasses several important downstream tasks: point cloud semantic segmentation, instance segmentation, and object detection, becomes increasingly important with the wide deployment of 3D sensors, such as LiDAR and RGB-D cameras [1, 2, 3, 4, 5, 6].
 Point clouds are raw sensor data obtained from 3D sensors and the most simple and common 3D data representation for understanding 3D scenes of robot navigation, robot grasping, and manipulation tasks.
@@
 Large-scale 3D scenes are very laborious to label, which also makes it very hard for deep network models to perform well with very limited annotations.
+Figure 2: The pre-training paradigm of our proposed WS3D++.
+We propose the hierarchical global to local feature alignments to establish the hierarchical vision-language aligned feature representations during the pre-training.
+This proposed paradigm helps to learn more powerful visual-linguistic aligned feature representation during the pre-training stage.
+We have further shown the final visualizations and comparisons with CLIP text presentations ranging from both the global view level to the local object category level.
+The results have further demonstrate the vision-language aligned feature representation for 3D scene parsing.
 Close-set assumption: One of the major bottlenecks in scaling up visual perception systems is the poor generalization capacity while encountered with diverse novel semantic classes or severe domain shifts.
@@
 These limitations severely restrict the scalability of VLM to handle diverse unseen 3D scenes containing diverse novel 3D object categories.
+Figure 3: The feature matching visualization of our proposed WS3D++.
+We propose hierarchical global to local feature alignments to establish hierarchical vision-language aligned feature representations during pre-training from both the global view level to the local object level.
+This kind of paradigm helps to learn more powerful visual-linguistic matched representations ranging from both the global view-level to the local object category-level.
+In the above figure, we have shown the matching at the global view on the left and the matching at local object level on the right.
+It can be demonstrated that our proposed approach can establish matched feature representation at both the global room feature level and the local object feature level.
+Figure 4: The fine-tuning paradigm of our proposed WS3D++.
+WS3D++ [4] consists of three proposed modules: 1.
+The unsupervised region-level energy-based optimization guided by boundary labels; 2.
+The unsupervised multi-stage region-level contrastive learning with high confidence; 3.
+The supervised region-level semantic contrastive learning with labeled data.
+The backbone network adopts encoder-decoder structures.
+The weights of the backbone network are shared in the supervised and unsupervised branches.
+Integrated with the proposed pre-training paradigm illustrated in Figure 2, by our proposed hierarchical feature aligned pre-training and regional fine-tuning, more effective label-efficient learning as well as open-vocabulary learning is realized.
 Reliance on large-scale labeled data.
@@
 And how to leverage boundary information has been explored in 2D but rarely explored in 3D data efficient learning [28].
-Second, the computation costs are both very high when applying point-level contrastive learning or point-level energy-based segmentation in a dense point cloud scene for every point pair [31].
+Second, the computation costs are both very high when applying point-level contrastive learning or point-level energy-based segmentation in a dense point cloud scene for every point pair [29, 30].
 Furthermore, large-scale point cloud scenes even contain billions of points, making point-level contrastive learning intractable in computational costs.
 Third, the existing unsupervised contrastive learning-based pre-training for point clouds [25, 20, 31, 32] only considers geometrically registered point/voxel pairs as the positive samples, while it does not explicitly consider explicit regional information, let alone the hierarchical alignments.
-Figure 1: Overall illustrative diagram of our proposed WS3D++.
-We integrate language-3D feature associated pre-training and data-efficient fine-tuning to achieve effective data-efficient and open-vocabulary 3D scene understanding for 3D scenes.
-Driven by the above motivations in terms of both generalization capacity and data efficiency, we propose an effective two-stage framework, involving unsupervised hierarchical vision-language pre-training and label-efficient fine-tuning to enable more label/data-efficient 3D scene understanding.
+Driven by the above motivations in terms of both generalization capacity and data efficiency, we propose an effective two-stage framework, involving unsupervised hierarchical vision-language pre-training and label-efficient fine-tuning to boost the label/data-efficiency in 3D scene understanding.
 As shown in Fig.
@@
 We also propose an effective knowledge distillation strategy that distills the informative visual-language-aligned representation of the image encoder in CLIP [7] to the 3D backbone network.
+As is demonstrated by our extensive experiments, the open-vocabulary scene parsing performance can be also enhanced.
 During the fine-tuning stage, as shown in Fig.
@@
 First, we propose a generalized pre-training approach for data-efficient learning, which establishes accurate alignments between language and 3D point cloud in terms of both object-level and scene-level semantics in a hierarchical manner.
-Second, we propose to leverage rendering techniques which make explicit associations between image and point cloud to facilitate 2D-to-3D matching and subsequent language-to-3D matching.
+Second, we propose to leverage the rendering technique that makes explicit associations between image and point cloud to facilitate 2D-to-3D matching and subsequent language-to-3D matching.
 Third, we visualized the language-queried activation maps directly on the 3D scenes, which demonstrates that the proposed approach learns better visual-linguistic alignment between the language descriptions and the visual object-level information.
@@
   1. (1)
-    During the pre-training stage, we first design an effective approach which distills rich knowledge from the large-scale vision-language model into the 3D point cloud modality.
+    During the pre-training stage, we first propose an effective design which distills rich knowledge from the large-scale vision-language model into the 3D point cloud modality.
 Specifically, we propose leveraging rendering to obtain explicit scene-level and object-level 2D-3D feature associations, establishing a more accurate vision-language association hierarchically than the original CLIP encoder.
+We have demonstrated by extensive experiments that our proposed approach can realize superior compatibility with prevailing weakly supervised approaches.
   2. (2)
-    During the pre-training stage, we further propose a word-to-3D matching approach to establish the well-aligned language-3D associated feature representations at both the scene level and the object level, which largely facilitates the subsequent effective contrastive learning with the mostly matched visual-linguistic contrastive pairs.
+    During the pre-training stage, we first propose a global scene-to-sentence matching and then propose a local object-to-word matching approach, respectively, to establish the well-aligned vision-language feature representations at both the scene level and the object level, which largely facilitates the subsequent effective contrastive learning with the mostly matched visual-language contrastive pairs.
 The proposed designs have both enhanced the data-efficient learning and the knowledge-transfer capacity of the model, as demonstrated by our extensive experiments on both the 3D object detection and the 3D semantic/instance segmentation tasks.
@@
 The energy-based loss and the contrastive loss are jointly optimized for pre-training the backbone network in a complementary manner, which take full advantage of the unlabeled data.
+    Figure 5: Qualitative semantic segmentation results of the proposed WS3D++ on SemanticKITTI validation set the with the 5% labeling percentage, compared with the fully supervised state-of-the-art Cylinder3D [34], and BAAF-Net [35] with the diverse semantics indicated by different colors.
+The red circles highlight the final performance difference between diverse comparative approaches.
   4. (4)
@@
 Our endeavor is orthogonal to the 3D backbone network designs and thus can be seamlessly integrated with the prevailing 3D point cloud detection or segmentation models.
-Our comprehensive results provide solid baselines for future research in data-efficient 3D scene understanding.
+Our comprehensive results provide solid baselines for future researches in the data-efficient 3D scene understanding.
+Figure 6: The captured 2D and 3D region proposals.
+It is demonstrated qualitatively clearly that more precise object proposals are captured by proposed united 2D/3D proposal generation approach.
+It can be demonstrated that clear superior regional proposal generation performance can be well guaranteed.
+Figure 7: The final ScanNet [18] object detection results and performance. The comparative differences in detection predictions are highlighted in the rectangles.
 2 Related Work
@@
 Deep-network-based approaches are widely adopted for point cloud understanding.
-Fully supervised approaches can be roughly categorized into voxel-based [36, 37, 38, 39, 40], projection-based [41, 42, 43, 44, 45, 46], and point-based approaches [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58].
+Fully supervised approaches can be roughly categorized into voxel-based [37, 38, 39, 40, 41], projection-based [42, 43, 44, 45, 46, 47], and point-based approaches [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58].
 The voxel-based approaches [59, 60, 61] which are built upon SparseConv [59] and voxelize the point cloud for efficient processing have achieved remarkable performance in 3D scene parsing.
 Therefore, we use SparseConv [59] as our backbone architecture for downstream semantic understanding tasks because of its high performance in inferring 3D semantics.
-Also, the LiDAR-based approaches are of significance to many industrial applications such as inspections [62, 63, 64] and robot localization [65, 66, 67, 68, 69], and large-scale robotic scene parsing [57, 70, 71], as well as robotic control applications [72, 73, 74, 75, 76, 77], etc.
 Pre-training for 3D Representation Learning.
@@
 Therefore, the designed pre-training approaches need to be very simple and lightweight, thus will make it easier to directly transfer the model to large-scale point clouds.
-Recently, with the development of large-scale vision-language models such as CLIP [9] and Flamingo [10], we can largely benefit the recognition capacity of 3D scene understanding models by distilling knowledge from large-scale vision-language models.
+Recently, with the development of large-scale vision-language models such as CLIP [7] and Flamingo [8], we can largely benefit the recognition capacity of 3D scene understanding models by distilling informative knowledge from large-scale vision-language models.
 For example, some pioneer works such as PointCLIP [66, 67] and ULIP [68] successfully transfer the knowledge from vision-language models to boost the downstream 3D shape classification tasks.
@@
 Compared with them, our proposed simple but effective framework can be both applicable to data-efficient learning and open-vocabulary scene understanding.
-Figure 2: The pre-training paradigm of our proposed WS3D++.
-We propose hierarchical feature alignments to establish hierarchical vision-language aligned feature representations during pre-training.
-This kind of paradigm helps learn more powerful visual-linguistic matched representations ranging from both the global view-level to the local object category-level.
 Label/Data-Efficient Learning for 3D.
-Recent studies have produced many elaborately designed backbone networks for 3D semantic/instance segmentation [88, 89, 45, 90, 91], as well as for 3D object detection [92, 93, 94, 95].
+Recent studies have produced many elaborately designed backbone networks for 3D semantic/instance segmentation [72, 35, 46, 73, 74], as well as for 3D object detection [75, 76, 77].
 However, they rely on full supervision.
@@
 More approaches [86] have been proposed recently, which utilize class prototypes and masked point cloud modeling [87, 88, 65] to learn informative representations for downstream 3D scene understanding.
+Conceptfusion [89] has also been proposed to alleviate the labelling burden via rendering with the proposed CLIP-driven queryable 3D point cloud maps.
 To sum up, although approaches have been proposed to alleviate the data efficiency problem, the models for weakly supervised learning lack the capacity to recognize novel categories beyond the labeled training set.
-Our framework tackles both open-set and data-efficient learning problems and is widely applicable to diverse 3D scene understanding downstream tasks.
+Our framework tackles open-set and data-efficient learning problems and is widely applicable to diverse 3D scene understanding downstream tasks.
+TABLE I: Comparisons of the open-vocabulary learning performance on ScanNet.
+It can be demonstrated that our proposed approach provides very superior open-world recognition performance compared with the diverse SOTAs.
+The results are given as hIoU / mIoUB / mIoUN, respectively.
+We have compared intensively with diverse labeled ratios.
+We apply panoptic quality mean intersection over union (mIoU), which can be divided into segmentation quality and recognition quality as metrics for the instance segmentation task.
+These evaluation metrics are computed on base (B) and novel (N) categories, with the superscripts of B and N (e.g.
+mIoUB / mIoUN), respectively.
+Furthermore, we have also utilized the harmonic metric such as harmonic IoU (hIoU) as the major indicators for open-world tasks following popular zero-shot learning work to consider the partition between base and novel classes.
+Datasets Models 80% 100%
+B15/N4 B12/N7 B10/N9 B15/N4 B12/N7 B10/N9
+ScanNet [18] Baseline 39.7 / 48.8 / 38.9 37.9 / 44.9 / 42.6 31.9 / 38.8 / 33.9 43.6 / 50.7 / 39.6 38.8 / 48.9 / 45.8 33.9 / 42.9 / 38.8
+Merely the Pre-training Stage 55.9 / 52.9 / 48.8 46.9 / 48.8 / 46.9 43.9 / 46.7 / 43.8 69.6 / 67.8 / 59.8 55.8 / 50.9 / 48.8 53.6 / 47.9 / 46.8
+3DGenZ [90] 32.6 / 76.8 / 26.6 25.9 / 46.5 / 26.3 28.3 / 63.6 / 18.6 42.9 / 76.9 / 39.6 32.9 / 56.8 / 33.9 27.1 / 69.8 / 32.8
+3DGenZ [90] (Pr) 42.8 / 78.8 / 42.9 39.6 / 68.3 / 51.8 27.6 / 79.5 / 18.9 46.8 / 88.6 / 48.6 43.8 / 78.8 / 49.9 37.9 / 83.7 / 38.9
+3DTZSL [91] 17.6 / 45.8 / 6.1 23.8 / 36.6 / 12.8 13.8 / 67.5 / 15.7 42.9 / 49.6 / 16.7 27.8 / 38.9 / 15.9 16.9 / 69.8 / 18.8
+3DTZSL [91] (Pr) 49.6 / 58.7 / 48.9 38.8 / 50.6 / 38.6 33.9 / 69.8 / 33.9 59.6 / 69.8 / 59.6 48.9 / 59.6 / 48.7 38.6 / 73.9 / 39.6
+LSeg3D [92] 0.0 / 64.4 / 0.0 26.9 / 55.7 / 22.8 11.8 / 77.6 / 23.9 33.6 / 76.8 / 33.8 28.8 / 72.6 / 29.2 26.6 / 68.5 / 26.9
+LSeg3D [92] (Pr) 38.9 / 78.8 / 29.9 32.8 / 73.6 / 26.8 28.6 / 66.8 / 22.8 49.6 / 82.7 / 46.9 46.9 / 76.8 / 43.2 38.9 / 70.3 / 39.6
+PLA without caption [69] 39.7 / 73.3 / 28.0 31.5 / 72.3 / 25.6 28.7 / 68.6 / 23.5 45.7 / 72.9 / 36.9 31.9 / 77.8 / 26.8 29.8 / 79.8 / 28.3
+PLA without caption [69] (Pr) 53.9 / 82.6 / 45.8 48.6 / 80.6 / 40.8 42.6 / 75.5 / 37.8 56.7 / 84.3 / 52.9 53.9 / 83.8 / 45.3 52.1 / 80.8 / 45.1
+PLA [69] 69.2 / 70.6 / 66.9 60.9 / 68.9 / 59.9 58.2 / 78.6 / 73.8 69.3 / 73.3 / 72.2 62.7 / 73.6 / 61.3 65.6 / 81.8 / 76.8
+PLA (Pr) [69] 79.9 / 82.9 / 81.9 78.5 / 80.6 / 68.9 65.9 / 77.8 / 57.8 82.5 / 88.2 / 84.9 80.8 / 85.7 / 73.9 76.8 / 85.3 / 63.9
+WS3D++ (Ours) 72.3 / 69.1 / 73.3 68.6 / 65.6 / 68.6 73.2 / 71.9 / 71.6 77.9 / 76.6 / 78.6 73.8 / 72.8 / 75.8 71.9 / 72.8 / 73.9
+WS3D++ (Ours) (Pr) 75.8 / 73.8 / 77.5 78.6 / 75.3 / 78.9 71.8 / 85.6 / 79.5 86.7 / 83.7 / 85.8 82.7 / 80.9 / 83.6 78.9 / 88.9 / 83.8
+Datasets Models 10% 50%
+B15/N4 B12/N7 B10/N9 B15/N4 B12/N7 B10/N9
+ScanNet [18] Baseline 27.9 / 18.2 / 21.9 23.0 / 33.6 / 31.2 19.6 / 28.6 / 27.6 35.1 / 20.6 / 36.6 31.7 / 41.5 / 38.2 28.5 / 32.3 / 28.2
+Merely the Pre-training Stage 35.8 / 58.3 / 36.8 33.8 / 37.9 / 31.8 30.8 / 70.3 / 25.6 37.2 / 63.2 / 39.7 48.9 / 53.8 / 49.9 49.6 / 86.8 / 49.8
+3DGenZ [90] 17.3 / 52.3 / 9.7 15.3 / 28.8 / 12.5 11.2 / 62.5 / 5.2 18.6 / 52.8 / 11.6 17.8 / 33.6 / 12.9 11.8 / 72.3 / 6.3
+3DGenZ [90] (Pr) 28.5 / 72.9 / 28.2 23.6 / 38.9 / 23.6 15.6 / 68.9 / 12.9 29.7 / 75.6 / 29.7 29.5 / 45.3 / 45.7 23.3 / 43.9 / 11.9
+3DTZSL [91] 8.9 / 33.9 / 6.1 3.3 / 35.8 / 1.8 7.3 / 47.3 / 2.8 16.2 / 32.5 / 23.2 21.9 / 31.3 / 6.8 9.9 / 63.5 / 8.8
+3DTZSL [91] (Pr) 26.8 / 43.8 / 28.7 18.6 / 52.6 / 23.2 16.8 / 49.5 / 19.2 31.3 / 49.3 / 28.7 28.6 / 43.9 / 27.3 15.3 / 66.9 / 21.7
+LSeg3D [92] 0.0 / 55.6 / 0.0 0.6 / 43.6 / 0.3 1.2 / 56.1 / 0.8 19.9 / 65.3 / 13.6 12.9 / 56.8 / 8.7 9.9 / 58.5 / 8.7
+LSeg3D [92] (Pr) 28.8 / 65.3 / 23.8 26.7 / 62.8 / 16.7 23.7 / 62.8 / 22.9 32.8 / 76.7 / 28.7 28.9 / 73.9 / 26.9 26.8 / 73.2 / 25.6
+PLA without caption [69] 35.3 / 53.2 / 18.2 26.7 / 52.1 / 9.9 19.6 / 56.7 / 9.7 36.5 / 55.9 / 22.8 28.2 / 58.7 / 18.8 25.8 / 62.1 / 15.2
+PLA without caption [69] (Pr) 39.3 / 73.3 / 35.7 28.6 / 67.8 / 29.6 24.5 / 62.6 / 26.8 42.9 / 75.6 / 41.8 39.3 / 73.6 / 36.9 35.6 / 72.3 / 33.8
+PLA [69] 51.3 / 58.8 / 48.6 49.3 / 55.9 / 44.9 44.7 / 52.2 / 40.8 65.7 / 72.8 / 65.7 58.8 / 66.5 / 49.8 53.6 / 63.5 / 46.8
+PLA (Pr) [69] 61.8 / 67.6 / 73.9 58.7 / 59.3 / 56.8 68.8 / 82.9 / 47.6 72.5 / 82.8 / 79.9 71.8 / 81.9 / 61.8 70.3 / 83.8 / 49.8
+WS3D++ (Ours) 65.9 / 75.2 / 68.3 61.3 / 65.6 / 62.7 58.6 / 60.9 / 56.8 71.8 / 76.8 / 81.6 65.8 / 69.5 / 68.9 62.8 / 67.8 / 68.2
+WS3D++ (Ours) (Pr) 69.9 / 68.8 / 67.8 67.9 / 67.7 / 66.9 63.6 / 66.2 / 65.6 70.9 / 70.8 / 71.9 72.8 / 71.8 / 69.8 69.8 / 75.6 / 72.1
 3 Proposed Methodology
@@
 2 and fine-tuning stage illustrated in Fig.
-3.
+4.
 During the pre-training stage, we first propose the hierarchical contrastive learning strategy with the help from rendering for more accurate vision-language alignments at both the scene level and object level.
@@
 Finally, during the fine-tuning stage, we directly propose a weakly supervised approach, we perform fine-tuning with regional boundary awareness and region-aware instance discrimination, which significantly improve model discrimination capacities when the labeled data are rather scarce.
-Figure 3: The fine-tuning paradigm of our proposed WS3D++, i.e.
-WS3D.
-WS3D [4] consists of three modules: 1.
-The unsupervised region-level energy-based optimization guided by boundary labels; 2.
-The unsupervised multi-stage region-level contrastive learning with high confidence; 3.
-The supervised region-level semantic contrastive learning with labeled data.
-The backbone network adopts encoder-decoder structures.
-The weights of the backbone network are shared in the supervised and unsupervised branches.
-Integrated with the proposed pre-training paradigm illustrated in Figure 2, by our proposed hierarchical feature aligned pre-training and regional fine-tuning, more effective label-efficient learning as well as open-vocabulary learning is realized.
 3.1 Hierarchical Vision-Language Knowledge Associations and Distillations for Pre-training
@@
 In terms of mesh-to-image transformation, we leverage the rendering pipeline including the vertex transformation, projection, and rasterization, as well as shading.
-We directly use the rendering library OpenGL [110] to render images from the meshes.
+We directly use the rendering library OpenGL [96] to render images from meshes.
 The process involves projecting the 3D vertices onto a 2D image plane based on camera parameters and applying shading and lighting calculations of 3D meshes to determine the specific color of each pixel.
-By this simple design, the world-to-camera extrinsic matrix $T_{e}$ containing both rotation and translation information between the 2D pixels and 3D points can be easily obtained.
+By our simple rendering design, the world-to-camera extrinsic transformation matrix $T_{e}$ containing both rotation and translation information between the 2D pixels and 3D points can be easily obtained.
 2D to 3D Alignment.
@@
 Moreover, we can find an explicit association between each element of the textual feature $\textbf{F}_{T}$ and the 3D feature $\textbf{F}_{3D}$ while passing through the backbone network.
+Figure 8: The object detection comparisons on KITTI [98] validation set.
+It can be demonstrated that our proposed WS3D++ can provide very accurate bounding box predictions qualitatively compared with the previous state-of-the-art merely using 2D bounding boxes, which demonstrates the effectiveness of using 3D regional as well as 3D object-level information in facilitating effective feature representation learning.
+Figure 9: Segmentation result comparisons with CLIP prompts for the indoor ScanNet benchmark.
+It can be seen from our results that clear object-level vision-language matched information can be captured with the designed visual prompts.
+It reveals the effectiveness of our designed hierarchical visual linguistic feature-aligned representation learning approaches.
 United 2D and 3D Proposal Generation.
 As shown in Fig.
-6, according to our experiments, we found that it is difficult to directly find the object-level information merely based on the 2D proposals due to the great information loss while conducting rendering, the proposals provided by the 2D region proposal network (RPN) [112] can not effectively capture the 3D object information within the scene.
-On the other hand, merely relying on the 3D proposal provided by the 3D RPN [113] still can not guarantee accurate proposal generation for the fact that some objects are too adjacent in their geometry.
-To this end, we propose to leverage the union of 2D and 3D RPN to capture intact and all-inclusive proposals within the 3D scenes.
+6, according to our experiments, we found that it is difficult to directly find the object-level information based on the 2D proposals merely due to the large information loss while rendering, the proposals provided by the 2D region proposal network (RPN) [99] can not effectively capture the 3D object information within the holistic scene.
+On the other hand, merely relying on the 3D proposals provided by the 3D RPN [100] still can not guarantee accurate proposal generation for the fact that some objects are too adjacent in their geometry.
+To this end, we propose to leverage the union of 2D and 3D RPN to capture intact and all-inclusive holistic proposals within the 3D scenes.
 Denote the region proposals as $R_{2D}$ and $R_{3D}$ respectively, the final holistic proposal generation $R_{H}$ is formulated as $R_{H}=R_{2D}\cup R_{3D}$.
-According to our experiments, this simple design can considerably boost the performance for the fact that the objects are more clustered and closely distributed within the indoor scenes.
+According to our experiments, this simple design can considerably boost the performance for the fact that the objects are more clustered and closely distributed within the indoor 3D scenes.
 Taking the union of 2D and 3D object proposals into consideration also guarantees that the optimization merely considers the regions where the object really exists and prevents the models from taking the pure background into consideration during the optimizations.
-Scene-level Sentence-3D Matching.
-We perform image-language alignment at both the object level and the scene level to pre-train the backbone network, which captures very accurate language-2D association for rendered images.
-The alignment at the object level allows the capturing of the fine-grained object-to-semantic concept matching, while the alignment at the scene level captures both holistic scene-level information and the dependent relations between the objects and holistic 3D scenes.
-For example, the bed exists in the bedroom while the sofa and TV lie in the living room with a clear owner-member relationship.
-To capture the object proposal information from the rendered images, we utilize the region proposal network (RPN) [112] to obtain the class-agnostic regions.
-For the text descriptor within the region, we use three most similar noun phase extract by it given by CLIP [9], which captures the scene level and the object level image patches with the most relevant texture description.
-Then we pass through both the original scene-level image as well as the obtained region-level image into the frozen CLIP visual encoder to obtain the scene-level visual feature $\textbf{I}_{\textbf{g}}$ as well as the object-level visual feature $\textbf{I}_{\textbf{l}}$, respectively.
-As shown in Figure 2, in order not to influence the well-learned feature representations provided by the CLIP encoder, we freeze the backbone and merely add a linear projection layer at the end of the network to facilitate the adaptive hierarchical feature-aligned pre-training.
-Global and Regional Word-3D Matching.
-At the next step, we perform regional word-to-3D matching.
-We use the CLIP image encoder to generate a set of local feature embeddings $\textbf{F}^{l}_{3D}$ for the rendered proposal-level image patches, and the set of global embedding $\textbf{F}^{g}_{3D}$ for the whole scene.
-While the text encoder extracts the local textual embeddings $\textbf{F}^{l}_{T}\in\mathbb{R}^{H\times D}$ and the global textual embeddings $\textbf{F}^{g}_{T}$.
-The operation can be interpreted as a bi-directional operation, which means that for each region proposal image patch, we find the textual concept that matches best with the semantics of the region.
-And for each textual concept, we find the multiple regions that correspond with it.
-Next, we calculate the cosine similarity $\mathcal{G}_{sim}$ between the 3D visually encoded features and the textual features:
-$\textbf{S}(\textbf{I},\textbf{T})=\mathcal{G}_{sim}(\textbf{F}^{l}_{3D},\textbf{F}^{l}_{T})+\mathcal{G}_{sim}(\textbf{F}^{g}_{3D},\textbf{F}^{g}_{T})$ (1)
+According to our previous experiments, although the average precision and recall can merely achieve 38.7/45.9% and 47.6/56.8% for 2D and 3D object proposal generation, respectively, the 3D scene parsing performance can still be well-maintained.
+By combining the 2D proposal with the 3D using the union operation while omitting the duplicated 2D proposals, the object proposal generation precision and recall can achieve a very high score of 65.2% and 76.8%, respectively, which demonstrates the superior effectiveness of them in generating region proposals.
+In the next stage, we perform vision-language matched contrastive candidate selection at both the object level and the scene level.
+The object-level vision-language matching makes alignment between the individual object and the descriptive word.
+While the scene-level vision-language matching make alignment between the whole scene with the descriptive sentence.
+In this way, the vision-language matched candidate can be selected for effective constrastive optimizations taking both global and local feature representation into account.
+Although the accuracy of proposal generation is not that high, our proposed approach can guarantee the final scene parsing performance as reported and demonstrated in Tables 1-8.
+The reason behind the phenomenon can be explained by that as long as the region proposal can be found, our framework can leverage the well-aligned representation in the vision language model (CLIP) to boost the final 3D semantic scene parsing performance, both in the open-vocabulary and closed-set scenarios.
+As validated by the previous experimental results and our qualitative validation in Figure 6, the quality of object proposal can be guaranteed, if the domain gap is not that large.
+Global 2D Scene-to-Sentence Matching.
+In the first place, we perform the global scene-to-sentence matching.
+For the rendered 2D image of the 3D scene, we utilize the GPT-4 to obtain the direct sentence-level description of the holistic scene.
+The sentence is given by the most similar description generated by GPT-4.
+Benefiting from the previous effective multi-view rendering designs, our proposed approach can have a holistic multi-view abstraction of the 3D environment.
+Denote the extract global scene-level visual feature as $V_{global}$ and denote the global text-level feature as $L_{global}$, then we can evaluate the alignment between visual and linguistic feature utilizing the similarity between $V_{global}$ and $L_{global}$.
+Local 2D Object-to-Word Matching.
+At the next step, we perform local object-to-word matching.
+In contrast to the global scene to scentence matching, here we conduct local object-to-word matching, which establishes and enhances vision-language association at more fine-grained object level and word level.
+The association is established at the object level for visual part and at the word level for language part.
+Specifically, we use the CLIP image encoder to generate a set of local visual feature embeddings $V_{l}ocal$ and utilize the text encoder to output a set of local word-level feature embeddings $L_{l}ocal$.
+In the next step, we calculate aggregated similarity using the cosine similarity $\mathcal{G}_{sim}$ between the 3D visually encoded features and the textual features:
+$\textbf{S}(\textbf{V},\textbf{L})=\mathcal{G}_{sim}(V^{2D}_{gobal},L^{2D}_{gobal})+\mathcal{G}_{sim}(V^{2D}_{local},L^{2D}_{local})$ (1)
 […]
-Contrastive Optimizations.
+2D-3D Visual Feature Distillation The main purpose is to conduct 2D-3D Visual Feature Distillation to obtain the aligned 2D-3D visual feature in the 2D-3D-language co-embedded feature space.
+We use the global 3D backbone to extract the 3D visual feature representation $V^{3D}_{global}$, and then utilize the $\mathcal{KL}$ divergence as the distillation loss to further distill the vision-language aligned informative knowledge from 2D feature space to 3D.
+This process can also be interpreted as the 2D-3D explicit feature alignment/distillation process.
+Compared with the mean square error loss such as the $\mathcal{L}_{1}$ or $\mathcal{L}_{2}$ losses, the $\mathcal{KL}$ divergence has improved regression capacity and ensured smoother gradient, which to some extent overcomes overfitting problems while distilling knowledge.
+The $\mathcal{KL}$ distillation loss $\mathcal{L}^{Dist}_{\mathcal{KL}}$ finally operates on the final two normalized 2D/3D vector for feature alignment:
+[…]
+Contrastive Language-Vision Optimizations.
 Note that when conducting contrastive learning, we regard textual features as anchors because textual descriptions are highly semantic and contain rich information, whereas the images contain too much low-level information and pixel-level details.
-Denote the $\textit{{F}}^{\;b,+}_{3D}$ and the $\textit{{F}}^{\;b,-}_{3D}$ as the positive and the negative features with respect to the anchor textual feature $\bm{F}_{T}^{a}$, respectively, the designed contrastive loss is formulated as follows:
-[…]
-$\mathcal{L}^{total}_{Ctr}=\mathcal{L}^{G}_{Ctr}+\mathcal{L}^{L}_{Ctr}$ (3)
-3D Language Distillation.
-We utilize the $\mathcal{KL}$ divergence as the distillation loss to further distill the knowledge from 2D feature space to 3D.
-Compared with the mean square error loss such as the $\mathcal{L}_{1}$ or $\mathcal{L}_{2}$ losses, the $\mathcal{KL}$ divergence has improved regression capacity and ensured smoother gradient [115], which to some extent overcomes overfitting problems while distilling knowledge.
-The distillation loss $\mathcal{L}^{Dist}_{\mathcal{KL}}$ is formulated as:
+Denote the $\textbf{V}^{3D,+}_{global}$ and the $\textbf{L}^{3D,-}_{global}$ as the positive and the negative feature with respect to the anchor textual feature $\bm{F}_{T}^{a}$, respectively, the designed contrastive discrimination loss is formulated as follows:
 […]
@@
 […]
-According to our extensive experiments, our simple pre-training approach provides well-aligned vision-language-3D co-embedding.
-It largely facilitates vision-language-associated knowledge transfers from 2D to 3D, boosting both the label efficiency and the final recognition capacity of novel categories.
+According to our extensive experiments, our simple pre-training approach provides the well-aligned vision-language-3D aggregated co-embedding.
+It considerably facilitates vision-language-associated knowledge transfers from 2D to 3D, boosting both the final label efficiency and the final recognition capacity of unseen novel categories.
 3.2 Region-Aware Fine-tuning
 During the fine-tuning stage, our proposed framework consists of three subparts for the network optimization: 1.
-Unsupervised energy-based loss guided by boundary awareness and highly confident network predictions for unlabeled data, which is discussed in 3.2.1; 2.
-Unsupervised multi-stage region-level contrastive learning with highly confident predictions for unlabeled data, which is discussed in 3.2.2.
+Unsupervised energy-based loss guided by boundary awareness and highly confident network predictions for unlabeled data, which is discussed in our original ECCV conference work [4]; 2.
+Unsupervised multi-stage region-level contrastive learning with highly confident predictions for unlabeled data.
 3.
-Supervised semantic contrastive learning for labeled data, which is discussed in 3.2.3.
-The three modules above are integrated jointly into the optimization function for network training to accomplish the final downstream detection or segmentation tasks with a very limited labeled data and a large amount of unlabeled data.
-3.2.1 Unsupervised Region-level Boundary Awareness
-Energy-function-based conditional random field segmentation is proposed in [29] and has been widely applied.
-However, it works in a fully supervised manner and does not explicitly consider the semantic boundary information, which is a great indication of semantic partitions in 3D scenes.
-In this Subsection, we develop a boundary-aware energy-based loss for unsupervised learning.
-As shown in Fig.
-1, to obtain robust boundaries for unlabeled 3D points, we first perform 3D over-segmentation [34] and also extract boundary points using an off-the-shelf semantic boundary prediction network, which is both subsequently used as the conditions to define the boundary regions for 3D points.
-Then, we propose a region-level energy-based loss based on obtained boundary region labels.
-Point Cloud Over-Segmentation.
-To obtain boundary regions, and facilitate subsequent region-level affinity computation and region-level contrastive learning, we first perform a region-level coarse clustering-based on point cloud over-segmentation.
-The previous method depended on the region growing [34] for over-segmentation, which relied heavily on the accurate normal estimation and was easily influenced by noises.
-In our work, we choose to use normal, curvature, and point feature histogram (PFH) [116] simultaneously to provide initial over-segmentation.
-The detailed procedures are given in the Appendix.
-And the over-segmentation results are shown in Fig.
-1 and 5.
-Denote original point clouds as $P_{in}$.
-After over-segmentation, they are partitioned into $Q$ subregions $S=\{s_{1},s_{2},...,s_{q}\}$, where ${s_{i}}\cap{s_{k}}=\emptyset$ for any ${s_{i}}\neq{s_{k}}$ as shown in Fig.
-1 and 5.
-Boundary Points Extraction.
-As shown in Fig.
-1, in addition to the over-segmentation result, we first extract the semantic boundary points to further identify boundary regions.
-The semantic boundary often indicates the distinguishment between various semantic classes.
-We extract semantic boundary points by JSENet [117], as shown in Fig.
-3.
-Note that the JSENet is trained with diverse ratios of labeled data in weakly supervised manners.
-As for training, we first define semantic boundary points from the limited labeled scenes as ground truth.
-With the definition of the ground truth boundary points, we design the loss following JSENet except substituting the binary cross entropy loss $L_{bce}$ with the focal loss $L_{foc}$ [118] to tackle large class imbalances between the boundary points and non-boundary points.
-$L_{foc}$ can be formulated as follows:
-$L_{foc}=-\frac{1}{N_{i}}\sum_{i=1}^{N_{i}}{(1-b^{gt}_{i})}^{\alpha}log(b_{i})+{(b^{gt}_{i})^{\alpha}}log(1-b_{i}),\vspace{-0.58 mm}$ (6)
-where $b_{i}$ denotes the binary predicted boundary map and $b^{gt}_{i}$ denotes the ground truth boundary map.
-$N_{i}$ is the total number of input points for training.
-We select $\alpha$=2 based on the original design [118].
-After its convergence, we apply the trained network to the remaining unlabeled scenes to obtain their boundary points.
-Examples of predicted boundary points of ScanNet [20] are shown in Fig.
-5, which clearly reveal distinctions between diverse semantic classes.
-Figure 4: Qualitative semantic segmentation results of the proposed WS3D++ on SemanticKITTI validation set the with the 5% labeling percentage, compared with the fully supervised state-of-the-art Cylinder3D [119], and BAAF-Net [89] with the semantics indicated by different colors.
-The red circles highlight the final performance difference between diverse approaches.
-Boundary Labels.
-After extracting semantic boundary points, we utilize them as labels of discrimination between diverse semantic categories.
-As shown in Fig.
-1, denote the $j_{th}$ adjacent regions of the center region $s_{i}$ as $s_{i,j}$.
-The adjacent region query is realized by fast Octree-based K-nearest neighbor search [120].
-Then, we determine the two adjacent regions as boundary regions if both $s_{i}$ and $s_{i,j}$ contain boundary points.
-The label for boundary region $w_{i,j}$ is designed as:
-$w_{i,j}=\begin{cases}1&\textit{if $s_{i}$, $s_{i,j}$ both contain boundary points;}\\ 0&\textit{otherwise.}\end{cases}$ (7)
-The label $w_{i,j}$ denotes semantic boundaries of adjacent regions, which is then used to guide the optimization of the energy function for the downstream segmentation task.
-Energy Loss Guided by Boundary Labels.
-As shown in Fig.
-3, we first perform data augmentation (detailed in the Appendix) for the input point clouds $\textbf{P}_{in}$ to obtain two transformed point clouds $\textbf{P}_{in,1}$ and $\textbf{P}_{in,2}\in\mathbb{R}^{N\times C_{in}}$, where N is the numbers of points.
-$C_{in}$ and $C_{out}$ are the numbers of input and output feature channels, respectively.
-Utilizing the backbone network SparseConv [59], we can obtain point cloud predictions $\textbf{P}_{out,1}$ and $\textbf{P}_{out,2}$.
-Then applying region-level max pooling on the same subregions, we obtain the predicted classes $\textbf{P}_{1}$, and $\textbf{P}_{2}\in\mathbb{R}^{M\times C_{out}}$ of the specific subregions.
-$\textbf{P}=\{p(s_{1}),p(s_{2}),...p(s_{i}),...,p(s_{R})\}$, where $R$ is the total number of regions obtained by over-segmentation.
-Denote the prediction of the $j_{th}$ neighboring region of the center region $s_{i}$ as $p(s_{i,j})$.
-Taking the unary network prediction and pairwise affinity between the neighbouring region into account, inspired by conditional random field (CRF) in DeepLab [29], we formulate the optimization energy function $E_{sum}$ as follows for label-efficient learning:
-$E_{sum}=\sum_{i}E_{i}(s_{i})+\sum_{i<j}^{adjacent}E_{i,j}(s_{i},s_{i,j}).$ (8)
-The first unary network prediction item $E_{i}(s_{i})$ is the entropy regularization term.
-It encourages region-level prediction with high confidence, which also facilitates the contrastive learning introduced subsequently, which is formulated as:
-$E_{i}(s_{i})=-log\,p(s_{i}).$ (9)
-We propose the pairwise affinity term of $E_{sum}$ as:
-$\begin{split}E_{i,j}(s_{i},s_{i,j})=H_{i,j}w_{i,j}[\epsilon-\|p(s_{i})-p(s_{i,j})\|]^{2}_{+}\\ +H_{i,j}(1-w_{i,j})\|p(s_{i})-p(s_{i,j})\|^{2}.\end{split}$ (10)
-$H_{i,j}$ is designed as the confidence indicator.
-$H_{i,j}=1$, if the probabilities which produce $p(s_{i})$ and $p(s_{i,j})$ are both larger than a threshold $\gamma$.
-Otherwise, it equals 00.
-$\epsilon$ can be any value in the range of $(0,1)$.
-$p(s_{i})$ and $p(s_{i,j})$ are the semantic predictions for the $i_{th}$ center region and the $j_{th}$ neighbouring region, respectively.
-And $[x]_{+}$ is the maximum function $max(0,x)$.
-For adjacent boundary regions, we encourage their confident semantic or instance predictions to be different (i.e.
-larger $\|p(s_{i})-p(s_{i,j})\|$); while for the non-boundary adjacent regions, we force their semantic predictions to be the same (i.e.
-$\|p(s_{i})-p(s_{i,j})\|=0$).
-Different from the traditional energy function in DeepLab[29], which used handcrafted features to compare similarities, we propose to use the learned boundary region labels to guide the network’s confident max-pooled region-level predictions.
-Therefore, the proposed boundary-aware energy function better encourages semantic separations at boundaries.
-Furthermore, we only consider pairwise affinity between adjacent regions instead of all pixel pairs, which greatly reduces computation costs, and avoids noises induced by distant unrelated pairs in the meanwhile.
-Figure 5: Qualitative instance segmentation results of proposed WS3D on ScanNet [20] with fully supervised state-of-the-art and the previous state-of-the-art GPC [121] in the 5% labeling ratio, compared with fully supervised arts.
-Different instances are indicated by different colors.
-And the intermediate over-segmentation results for obtaining regions and boundary predictions, with the boundary points indicated by the color blue.
-3.2.2 Region-level Instance Discrimination
-After applying the entropy regularization term $E_{i}(s_{i})$, we can obtain predictions at the region level with high confidence.
-Note that confident region-level predictions further improve the latent feature discrimination capacity of the network, which makes contrastive learning in the latent space feasible.
-Therefore, we further propose a multi-stage region-level contrastive learning for unlabeled data.
-Compared to previous work using only contrastive learning with low-level geometric registrations [27], our work unleashes potentials of contrastive learning with instance discrimination to enhance versatile feature learning in latent space.
-The key to successful semantic/instance segmentation is maintaining discriminative feature representations at different stages of the backbone network [122].
-Inspired by the pyramidal feature network [122, 123], we propose a simple but effective multistage contrastive learning approach for point clouds in an unsupervised setting.
-As shown in Fig.
-3, given the augmented point clouds input $P_{in,1}$ and $P_{in,2}$, we feed them into the backbone encoder.
-We add five additional MLP heads with region-level max-pooling to obtain region-level segmentation predictions at the $m_{th}$ backbone stage, denoted as $\textbf{f}_{g,m}$ and $\textbf{f}_{h,m}$, respectively (five stages in our case, denote $M$ as the total network stages.
-i.e., $M$ = 5).
-After we apply the MLP heads to the extracted features at different stages, we can obtain the hierarchical feature embeddings of the fine-tuned downstream 3D scene understanding network.
-Unlike existing pixel-level [124], or point-level [124, 125] for point cloud contrastive learning, our proposed contrastive learning performs at the region level.
-Therefore, it will help to obtain more holistic regional information.
-The obtained regional information will be complementary to the point-level information.
-The region-level semantic contrastive loss is formulated as:
-$L^{m}_{contrast}=-\frac{1}{\textbf{S}_{p}}\sum_{(a,b)\in S_{p}}log\frac{H_{i,j}exp(\textbf{f}_{g,m}^{a}\cdot\textbf{f}_{h,m}^{+}/\tau)}{\sum_{(\cdot,c)\in S_{p}}H_{i,j}exp(\textbf{f}_{g,m}^{a}\cdot\textbf{f}_{h,m}^{-}/\tau))},$ (11)
-where $(a,b)\in\textbf{S}_{p}$ are latent confident predicted positive region pairs, and $(a,c)\in\textbf{S}_{p}$ are latent negative region pairs.
-As mentioned before, $H_{i,j}$ is designed to eliminate contrastive learning candidates with low confidence degrees.
-Reliable region-level contrastive learning is only applied to confident predictions by the network.
-After applying contrastive losses for instance discrimination, semantically similar features with high confidence levels in the latent space will become more adjacent, whilst semantically distinct ones will be separated, which will be shown in Fig.
-13.
-Note that although a recent work GPC [121] has proposed methods to perform contrastive learning on the point clouds in a SSL manner, our work is different from their method in two aspects.
-Firstly, our contrastive learning is conducted at region level while GPC conducts contrastive learning at the point level.
-Secondly, GPC focuses on the selection of the positive and negative point-set samples to perform contrastive learning in a pseudo-label supervised manner on two different 3D scene samples, while we focus on the unsupervised contrastive learning which disentangles different feature representations in the latent space on the same augmented 3D scene sample, guided by confident network predictions, and it learns hirachical regional object information.
-The final proposed multi-stage contrastive learning loss is formulated as the sum of losses at every network stage:
-$L_{multi}=\sum_{m=1}^{M}L_{contrast}^{m},\vspace{-0.1mm}$ (12)
-After applying the multi-stage contrastive loss $L_{multi}$, the output at each stage of the network will provide more distinctive representations to attain better performance.
-From our ablation experiments, the performances can be boosted by applying multi-stage contrastive loss.
-We apply the typical self-training-based approach, where the high-confidence predictions are retained as the pseudo label for the network training in the subsequent iteration.
-We adopt the typical cross-entropy loss for the self-training which is denoted as $L_{ST}$.
-Considering $E_{sum}$ (see Eq.
-8) for the two augmented scenes $P_{in,1}$, $P_{in,2}$, and $L_{multi}$ (see Eq.
-12), we formulate the overall loss $L_{unlabeled}$ for the WS3D training with unlabeled data: $L_{unlabeled}=L_{ST}+\alpha E_{sum,2}+\beta L_{multi}$.
-3.2.3 Supervised Learning for labeled data
-We also guide network optimizations by using supervision from the labeled data.
-As shown in Fig.
-3, we use the cross-entropy loss $L_{ce}$ to guide the supervised learning on the labeled data in the supervised branch.
-The loss term for the WS3D training with the labeled data is $L_{labeled}$.
-For the task of semantic segmentation, the loss for the labeled data $L_{labeled}$ includes the cross-entropy loss.
-For the backbone network selection, we follow previous work with the sparse convolutional network design for semantic [59] and instance segmentation [90], respectively.
-Our detection network is designed based on widely adopted VoteNet [126, 99].
-We also use Dice [127] loss to guarantee tighter aggregations of points within the same cluster, and strict geometric separations of points in diverse clusters [99].
-3.2.4 The Overall Optimization Loss Function
-Leveraging our proposed region-level energy-based loss and region-level contrastive learning, the network can make maximum use of the unlabeled data for better feature learning to boost performance in the fine-tuning stage.
-As shown in Fig.
-3, for semantic segmentation and instance segmentation, we fine-tune the network in an end-to-end manner for both supervised and unsupervised branches to make full use of labeled and unlabeled data.
-The weight value of the unsupervised loss is denoted as $\lambda_{u}$.
-The overall optimization function $L_{total}$ is formulated as follows, and we set $\lambda_{u}=2$ empirically:
-$L_{total}=L_{labeled}+\lambda_{u}L_{unlabeled}.$ (13)
-Figure 6: The captured 2D and 3D region proposals. It is demonstrated that more object proposals can be captured by our proposed combined unitive 2D/3D proposal generation approach.
-Figure 7: The final ScanNet [20] object detection results. Differences in predictions are highlighted in the rectangles.
-Figure 8: The object detection comparisons on KITTI [128] validation set.
-It can be demonstrated that our proposed WS3D++ provides very accurate bounding box predictions qualitatively compared with the previous state-of-the-art merely using 2D bounding boxes, which demonstrates the effectiveness of using 3D regional as well as 3D object-level information in facilitating effective feature representation learning.
-Figure 9: Segmentation result comparisons with CLIP prompts for the indoor ScanNet benchmark.
-It can be seen from our results that clear object-level information can be captured with the designed visual prompts, which demonstrates the effectiveness of our designed hierarchical feature-aligned representation learning approaches.
+Supervised semantic contrastive learning for labeled data.
+The three important modules above are integrated jointly into the optimization function for network training to accomplish the final downstream detection or segmentation tasks with a very limited labeled data, with all remaining data unlabeled.
+Figure 10: The segmentation result visualizations and comparisons with CLIP text encoder prompts for the outdoor KITTI benchmark.
+It can be demonstrated the final foreground object awareness can be clearly captured as compared with the previous SOTAs approach PLA [69].
+Meanwhile, as shown in the last three columns, we can provide clear segmentation for the corresponding visual objects based on the textual prompts.
+The results further demonstrate that vision-language aligned representations can be effectively and sufficiently learnt.
 4 Experiments
@@
 Detailed information on each dataset and training details are put into the Appendix.
+The pre-training is conducted on ScanNet training set for indoor benchmark and on Waymo for outdoor benchmark.
 Training Set Partition.
@@
 One of the sub-scenes belongs to the labeled data and the other sub-scene belongs to the unlabeled data.
-TABLE I: Comparison of semantic segmentation results with different labeling percentages on ScanNet validation set, S3DIS validation set (Area 5), and SemanticKITTI validation set (Sequence 08).
+TABLE II: Boundary prediction average precision (AP) diverse label ratios for outdoor SemanticKITTI (SKITTI) benchmark and indoor ScanNet and S3DIS benchmark.
+Left of ’/’: The Average Precision (AP) of Semantic Boundary Prediction with focal Loss $L_{foc}$ Right of ’/’: The average precision with the cross entropy loss $L_{bce}$.
+Labeling The APof Boundary Prediction (BP) WS3D++ Sem. Seg. mIOU%
+Percentage SKITTI (Outdoor) ScanNet (Indoor) S3DIS (Indoor) SKITTI (Outdoor) ScanNet (Indoor) S3DIS (Indoor)
+1% 56.7 / 61.6 55.8 / 59.1 56.2 / 61.7 51.6 58.8 55.8
+5% 53.9 / 60.8 58.9 / 63.9 59.6 / 69.2 59.3 67.9 64.7
+10% 55.7 / 61.5 62.2 / 65.7 62.6 / 66.9 68.7 76.7 68.2
+15% 57.6 / 67.9 65.2 / 68.2 62.9 / 67.8 75.2 76.2 68.8
+20% 59.7 / 66.5 66.3 / 69.8 65.2 / 68.7 72.9 79.8 75.8
+25% 62.6 / 65.6 67.8 / 70.3 65.8 / 67.8 74.7 77.1 75.5
+30% 64.6 / 70.2 71.6 / 73.8 68.9 / 72.8 78.7 76.3 75.9
+40% 68.6 / 72.8 73.2 / 75.3 71.2 / 73.5 75.6 81.6 77.8
+100% 83.7 / 88.7 76.9 / 82.6 76.8 / 83.7 83.9 88.8 82.9
+TABLE III: Comparison of semantic segmentation results with different labeling percentages on ScanNet validation set, S3DIS validation set (Area 5), and SemanticKITTI validation set (Sequence 08).
 ’Sup-only-GPC’ denotes GPC model trained with only labeled data.
@@
 We have shown the performance increase in the last row for each dataset, compared to merely trained models with labeled data (left value) and to the SOTAs GPC [103] (right value).
+The results in brackets (Pr) mean that the results after using our proposed pre-training approach to increase the final performance.
+It should be noted that the pre-training models for all weakly supervised approaches are the same with our proposed WS3D++ pre-training.
+It can be demonstrated that our proposed pre-training approach has a significant boost on the final scene parsing performance.
+Note that our subscript highlights the performance increment compared with the ”Sup-only-GPC”.
 Datasets Approaches Semantic Segmentation mIoU (%) on the Validation Set According to Supervision Level (%)
 1% 5% 10% 15% 20% 25% 30% 40% 100%
-ScanNet Sup-only-GPC 40.9 48.1 57.2 61.3 64.0 65.3 67.1 68.8 72.9
+ScanNet Sup-only-GPC–Baseline 40.9 48.1 57.2 61.3 64.0 65.3 67.1 68.8 72.9
+Sup-only-GPC (Pr) 44.8 56.9 59.8 64.9 66.7 69.9 66.8 63.8 70.8
 GPC [103] 46.6 54.8 60.5 63.3 66.7 67.5 68.9 71.3 74.0
+GPC (Pr)[103] 52.5 59.3 67.8 68.9 69.7 69.8 75.3 76.5 76.6
 Mix3D [104] 47.1 55.2 61.1 63.5 66.8 67.9 68.7 71.5 73.3
+Mix3D [104] (Pr) 52.7 58.9 65.1 67.9 70.2 70.6 71.3 73.8 78.6
 PointMixup [105] 44.5 55.2 69.6 63.7 66.3 66.9 68.8 70.2 73.8
+PointMixup  [105] (Pr) 46.8 59.8 75.5 68.1 69.8 73.7 72.8 75.3 78.6
 RDPL [106] 47.6 56.3 61.1 63.8 66.8 67.8 69.7 71.8 74.6
+RDPL [106] (Pr) 50.7 59.6 66.9 68.3 71.6 71.5 72.8 73.9 76.8
 Active-ST [107] 48.7 56.7 62.5 65.6 67.8 68.7 71.2 72.5 75.8
-WS3D [4] 49.9 +9.0 56.2 +8.1 62.2 +5.0 65.8 +4.5 68.5 +4.5 69.4 +4.1 70.3 +3.2 73.4 +4.6 76.9 +4.0
-WS3D++ 52.6 +11.7 59.1 +11.0 65.2 +8.0 66.9 +5.6 71.8 +6.6 71.1 +5.8 72.6 +5.5 75.3 +6.6 81.7 +8.8
-S3DIS Sup-only-GPC 36.3 45.0 52.9 53.8 59.9 60.3 61.2 62.6 66.4
+Active-ST [107] (Pr) 49.9 59.5 67.9 67.9 75.3 70.6 73.7 75.2 78.1
+WS3D++ 52.6 +11.7 59.1 +11.0 65.2 +8.0 66.9 +5.6 71.8 +6.6 72.0 +6.7 73.9 +5.5 75.3 +6.6 81.7 +8.8
+WS3D++ (Pr) 58.8 +17.9 67.9 +19.8 76.5 +19.5 78.6 +14.9 79.8 +15.8 80.1 +11.8 80.8 +11.6 83.8 +12.8 88.8 +15.9
+Merely Pre-training Stage 43.9 52.3 56.3 61.6 65.3 67.6 68.5 69.8 71.9
+Sup-only-GPC–Baseline 35.3 44.5 52.9 53.8 59.9 60.3 61.2 62.6 66.4
+S3DIS Sup-only-GPC (Pr) 48.6 51.7 60.8 61.9 70.8 72.8 69.7 68.8 73.6
 GPC [103] 38.2 53.0 57.7 60.2 63.5 63.9 64.9 65.0 68.8
+GPC [103] (Pr) 47.8 58.7 63.8 66.8 68.5 69.7 68.8 68.9 71.2
 Mix3D [104] 39.2 50.3 58.3 61.1 64.1 64.8 65.9 66.9 69.6
-WS3D [4] 45.3 +9.0 54.6 +9.6 59.3 +6.4 62.3 +7.0 65.7 +5.8 66.5 +6.8 67.2 +6.0 69.5 +6.9 72.9 +6.5
-WS3D++ 48.6 +12.3 57.7 +12.7 61.2 +8.3 66.9 +11.6 70.6 +10.7 71.1 +10.8 72.6 +11.4 75.3 +12.7 81.7 +15.3
-SemanticKITTI Sup-only-GPC 28.6 34.8 43.9 47.9 53.8 55.1 55.4 57.4 65.0
+Mix3D [104] (Pr) 47.2 57.9 63.2 68.2 69.5 70.6 69.8 69.8 73.9
+WS3D++ 48.6 +12.3 57.7 +12.7 61.2 +8.3 66.9 +11.6 70.6 +10.7 71.1 +10.8 72.6 +11.4 75.3 +12.7 82.0 +15.6
+WS3D++ (Pr) 55.8 +19.5 64.7 +19.7 68.2 +15.3 68.8 +15.0 75.8 +15.9 75.5 +15.2 75.9 +14.7 77.8 +15.2 82.9 +16.5
+Merely Pre-training Stage 45.2 52.8 58.7 63.8 68.6 69.8 71.2 72.6 73.9
+SemanticKITTI Sup-only-GPC–Baseline 28.6 34.8 43.9 47.9 53.8 55.1 55.4 57.4 65.0
+Sup-only-GPC (Pr) 36.3 41.8 50.3 55.6 57.6 63.8 60.8 59.8 67.8
 GPC [103] 34.7 41.8 49.9 53.1 58.8 59.1 59.4 59.9 65.8
+GPC [103] (Pr) 36.9 45.6 53.8 55.8 59.9 63.9 65.6 64.8 69.5
 LESS [108] 37.1 42.5 50.5 53.9 59.5 59.6 60.5 63.5 66.9
+LESS [108] (Pr) 39.6 46.7 53.9 56.8 63.9 65.7 65.3 68.7 68.8
 Mix3D [104] 37.7 42.9 50.8 54.1 59.9 60.9 61.8 61.3 68.8
-WS3D [4] 38.9 +10.3 43.7 +8.9 52.3 +8.4 55.5 +7.6 61.4 +7.6 61.8 +6.7 62.1 +6.7 63.2 +5.8 66.9 +1.9
-WS3D++ 46.8 +18.2 58.6 +13.8 55.2 +11.3 62.8 +14.9 65.9 +12.1 67.9 +12.8 68.6 +13.2 70.9 +13.5 76.8 +11.8
-Nuscene Sup-only-GPC 39.8 45.9 47.1 48.6 55.7 56.3 58.9 59.8 67.9
-PRKT [135] 48.0 49.3 51.8 53.5 56.7 59.6 63.8 66.7 70.1
-SLidR [134] 48.2 50.3 52.9 55.8 56.9 59.1 60.8 63.8 66.9
-PointContrast [27] 48.3 50.2 55.1 58.1 60.2 61.1 62.1 64.9 69.2
-CLIP2Scene [27] 56.3 56.8 58.7 61.2 63.6 64.1 64.5 64.8 65.1
-LESS [134] 49.1 42.5 50.5 53.9 59.5 59.6 60.5 63.5 66.9
-Mix3D [130] 37.7 42.9 50.8 54.1 59.3 60.2 61.8 61.3 68.8
-WS3D [4] 49.1 +9.3 51.6 +5.7 52.3 +5.2 55.5 +6.9 61.4 +6.7 61.8 +5.5 62.1 +3.2 63.9 +4.1 71.6 +3.7
-WS3D++ [4] 53.7 +13.9 56.7 +10.8 55.2 +11.3 62.8 +14.9 71.6 +15.9 67.9 +11.6 68.6 +9.7 71.9 +12.1 77.8 +9.9
+Mix3D [104] (Pr) 41.8 46.6 53.9 58.6 64.7 65.6 66.6 65.9 73.7
+WS3D++ 46.8 +18.2 48.6 +13.8 55.2 +11.3 62.8 +14.9 65.9 +12.1 67.9 +12.8 68.6 +13.2 70.9 +13.5 76.8 +11.8
+WS3D++ (Pr) 51.6 +33.0 59.3 +24.5 68.7 +24.8 69.8 +21.9 72.9 +19.1 74.7 +19.6 76.3 +20.9 76.6 +19.2 83.9 +18.9
+Merely Pre-training Stage 42.9 46.6 51.7 61.7 63.8 67.8 69.6 71.8 75.8
 Figure 11: The final 3D scene-level activations results based on the language query of WS3D++. Better zoom in for details.
-[…]
+Figure 12: t-SNE visualization in semantic segmentation of the proposed WS3D under the 5% labeling percentage on S3DIS validation set.
+The diverse feature embeddings are indicated by different colors and are normalized into $[-1,1]$ for better visualization.
+It has been validated that more discriminative features can be acquired for diverse semantic classes with our proposed unsupervised region-aware fine-tuning strategy (demonstrated by WS3D) and our proposed hierarchical vision-language knowledge associated and distilled pre-training (demonstrated by WS3D++).
+Implementation Details.
+For the task of semantic segmentation, we fine-tune the network for 500 epochs on a single NVIDIA 1080Ti GPU with a batch size of 16 during training.
+The initial learning rate is set to 1$\times$10-3 and is multiplied with 0.2 every 50 epochs.
+We implement it in PyTorch and optimize it with Adam optimizer [109].
+We set the hyperparameter $\gamma$ as 0.8 to ensure that merely highly confident prediction can be used for network optimization.
+$\epsilon$ is set to 0.5.
+We empirically choose $\alpha=\beta=1$, while $\lambda_{u}=0.1$.
+For instance segmentation, we train the network for 580 epochs on a single NVIDIA 1080Ti GPU with a batch size of 8 during training.
+The other settings are the same as the semantic segmentation task.
 4.3 Data-efficient 3D Semantic Segmentation
 Overall Experimental Results.
-For the semantic segmentation, we have tested WS3D on versatile indoor and outdoor benchmarks, including ScanNet [20], S3DIS [21], and SemanticKITTI [18].
+For the semantic segmentation, we have tested WS3D++ on versatile indoor and outdoor benchmarks, including ScanNet [18], S3DIS [19], and SemanticKITTI [16].
 We have done extensive experiments with limited labeled data, e.g., only {1%, 5%, 10%, 15%, 20%, 25%, 30%, 40%, 100%} data in training set are available as labeled data.
 The qualitative results are shown in Fig.
-4.
-In the meanwhile, the quantitative semantic segmentation performance is summarized in Table I.
+5.
+In the meanwhile, the quantitative semantic segmentation performance is summarized in Table III.
 As mentioned, we have used the voxel-based method SparseConv [59] as the backbone.
@@
 For example, the performance increase on SemanticKITTI is 10.3% for the 1% labeling percentage, 5.8% for the 40% labeling percentage, and 1.9% for the 100% labeling percentage.
-This can be possibly explained by the fact that for more unlabeled data, our proposed WS3D can extract more meaningful semantic information from the unlabeled data based on our boundary-guided energy-based loss and confidence-guided region-level contrastive learning design.
-In addition, compared with current SOTA GPC, our proposed WS3D also achieves consistently better results in semantic segmentation performance, especially when faced with very limited label circumstances (e.g.
+This can be possibly explained by the fact that for more unlabeled data, our proposed WS3D++ can extract more meaningful semantic information from the unlabeled data based on our boundary-guided energy-based loss and confidence-guided region-level contrastive learning design.
+In addition, compared with current SOTA GPC, our proposed WS3D++ also achieves consistently better results in semantic segmentation performance, especially when faced with very limited label circumstances (e.g.
 1% labeling points).
-In that case, WS3D outperforms GPC by 3.3%, 7.1%, and 4.2% for ScanNet, S3DIS, and SemanticKITTI, respectively.
+In that case, WS3D++ outperforms GPC by 3.3%, 7.1%, and 4.2% for ScanNet, S3DIS, and SemanticKITTI, respectively.
 Fig.
 5 shows that we can provide comparable performance compared with fully supervised SOTAs BAAF-Net [35] and Cylinder3D [34] on SemanticKITTI with 5% labels.
-As shown in Table I, the performance of our enhanced approach WS3D++ has remarkably increased performance compared with WS3D and previous SOTAs, demonstrating the effectiveness of our proposed vision-language knowledge-associated pre-training.
-Figure 11: Segmentation result visualizations and comparisons with CLIP text encoder prompts for the outdoor KITTI benchmark.
-It can be demonstrated the foreground object awareness can be clearly captured comared with the previous SOTAs PLA [85].
-Meanwhile, as shown in the last three columns, we can provide clear segmentation for the corresponding visual objects based on the textual prompts.
-TABLE II: Comparison of experimental results on 20% and fully labeled case for the task of inductive and transductive learning for our proposed WS3D++, respectively.
+As shown in Table III, the performance of our enhanced approach WS3D++ has remarkably increased performance compared with WS3D++ and previous SOTAs, demonstrating the effectiveness of our proposed vision-language knowledge-associated pre-training.
+We have reported the experimental results about semantic boundary prediction as demonstrated in Table 1 of our revised manuscript, it is demonstrated that our proposed approach can provide a very high-quality label at a very low-label regime, and the boundary prediction average precision can be kept at more than 50% (50.9% at least).
+The phenomenon has also been found at previous research including SQN, Lasermix, and our Weaklab3DNet [110].
+The reason behind is that: As point clouds are essentially samples of the 3D world, the distribution of points in a very close local neighborhood is comparatively homogeneous, revealing strong semantic similarity/homogeneity.
+Moreover, our proposed weakly supervised approach can be regarded as an amplification of those rather sparse supervision signals, which largely facilitates ultimate semantic boundary prediction.
+As we have demonstrated in Table I, the average precision (AP) of boundary prediction can still be maintained at a relatively high value (more than 50.9%).
+Our proposed approach can realize promising performance on outdoor benchmark SemanticKITTI and indoor benchmarks including S3DIS and ScanNet, which demonstrates the superior effectiveness of our proposed weakly supervised semantic boundary prediction.
+Figure 13: The WS3D++ corresponding detection and segmentation projections on six rendered 2D views through our proposed multi-view rendering approach.
+It can be demonstrated that our proposed rendering has established an explicit modality association between the final 2D views and 3D views.
+Also, by combining 2D and 3D region proposals, more complete and apparent object-level information can be clearly captured both from 2D views and 3D views.
+Best zoom-in for viewing.
+TABLE IV: Comparison of experimental results on 20% and fully labeled case for the task of inductive and transductive learning for our proposed WS3D++, respectively.
 In transductive learning, the test set is also utilized for network training.
@@
 Base Induct. Transduct. Base Induct. Transduct.
-ScanNetv2-WS3D 64.0 68.5 71.4 72.9 76.9 77.6
-ScanNetv2-WS3D++ 67.3 73.9 75.6 77.3 80.3 81.6
-S3DIS Area5 Val. 59.9 65.7 66.6 66.4 72.9 73.5
-S3DIS WS3D 61.3 67.6 68.6 68.2 74.8 75.8
-S3DIS WS3D++ 62.8 68.9 69.9 69.7 75.5 76.9
-Semantic KITTI Val. 53.8 61.4 64.5 65.0 66.9 68.2
-Semantic KITTI Val. WS3D++ 54.7 62.9 66.2 67.1 68.2 69.7
-Semantic KITTI Test. 55.7 62.5 63.6 65.4 68.1 71.3
-Semantic KITTI Test. WS3D++ 56.6 63.7 65.3 67.1 69.7 72.8
-TABLE III: Comparison of current State-of-the-art (SOTA) approaches in the limited reconstruction case for the 3D object detection tasks with different ratios of labeled data.
-The mean average precision (mAP) given by the mean $\pm$ standard deviation in three runs of diverse random splits are reported.
-Datasets Models 5% 10% 20%
-Induct. Transduct. Induct. Transduct. Induct. Transduct.
-SUN RGB-D [137, 66] Baseline 29.9 $\pm$ 1.5 33.5 $\pm$ 0.8 34.4$\pm$ 1.1 40.7$\pm$ 0.9 41.1 $\pm$ 0.3 47.5 $\pm$ 0.5
-SESS [138] 34.2 $\pm$ 2.0 38.1 $\pm$ 0.7 42.9 $\pm$ 0.8 45.3$\pm$ 0.9 47.9 $\pm$ 0.4 51.6 $\pm$ 0.3
-3D-IOUMatch [139] 39.1 $\pm$ 1.9 46.3 $\pm$ 0.7 45.5 $\pm$ 1.5 53.5 $\pm$ 0.3 49.7 $\pm$ 0.5 54.3 $\pm$ 0.8
-SPD [140] 38.5 $\pm$ 0.7 44.3 $\pm$ 0.8 46.0 $\pm$ 1.0 51.7 $\pm$ 0.9 49.6 $\pm$ 0.5 54.5 $\pm$ 0.9
-WS3D (Ours) 46.9 $\pm$ 0.8 49.5 $\pm$ 0.7 47.8 $\pm$ 0.8 55.7 $\pm$ 0.8 55.6 $\pm$ 0.7 57.9 $\pm$ 1.2
-WS3D-Open (Ours) 47.8 $\pm$ 0.8 50.3 $\pm$ 0.6 49.3 $\pm$ 0.7 58.2 $\pm$ 0.7 59.1 $\pm$ 0.8 60.3 $\pm$ 1.1
-WS3D++ (Ours) 52.7 $\pm$ 0.7 54.5 $\pm$ 0.8 53.7 $\pm$ 0.9 63.8 $\pm$ 0.8 65.8 $\pm$ 0.7 65.6 $\pm$ 1.1
-ScanNet-V2 [20] Baseline 27.9 $\pm$ 0.5 30.8 $\pm$ 1.5 31.1 $\pm$ 0.7 33.2 $\pm$ 0.5 41.6 $\pm$ 0.9 44.5 $\pm$ 1.2
-SESS [138] 32.2 $\pm$ 0.8 36.8 $\pm$ 1.1 39.7 $\pm$ 0.9 44.5 $\pm$ 0.5 47.9 $\pm$ 0.4 49.2 $\pm$ 0.7
-3D-IOUMatch [139] 40.0 $\pm$ 0.9 46.3 $\pm$ 0.7 47.2 $\pm$ 0.4 49.6 $\pm$ 0.6 52.8 $\pm$ 1.2 54.3 $\pm$ 0.8
-SPD [140] 41.5 $\pm$ 0.5 44.3 $\pm$ 0.8 43.2 $\pm$ 1.2 46.2 $\pm$ 0.5 51.9 $\pm$ 0.5 55.1 $\pm$ 0.9
-WS3D (Ours) 45.8 $\pm$ 0.5 49.2 $\pm$ 0.6 47.7 $\pm$ 0.9 53.9 $\pm$ 0.6 55.6 $\pm$ 0.7 58.2 $\pm$ 1.3
-WS3D-Open (Ours) 47.2 $\pm$ 0.8 50.8 $\pm$ 0.6 51.1 $\pm$ 0.9 59.2 $\pm$ 0.8 59.0 $\pm$ 0.8 61.5 $\pm$ 1.6
-WS3D++ (Ours) 55.3 $\pm$ 1.1 57.6 $\pm$ 1.2 57.9 $\pm$ 0.8 65.8 $\pm$ 1.2 64.6 $\pm$ 0.5 66.5 $\pm$ 1.1
-TABLE IV: Comparisons of the open vocabulary learning performance.
-It can be demonstrated that our proposed approach provides very superior open-world recognition performance compared with the diverse SOTAs.
-WS3D-Open abandons our feature-aligned contrastive pre-training with the contrastive loss term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [9] feature encoder with the knowledge distillation loss term $\mathcal{L}^{Dist}_{\mathcal{KL}}$.
-Datasets Models Few-shot settings
-B15/N4 B12/N7 B10/N9
-ScanNet [20] 3DGenZ [141] 20.6/56.0/12.6 19.8/35.5/13.3 12.0/63.6/6.6
-3DTZSL [142] 10.5/36.7/6.1 3.8/36.6/2.0 7.8/55.5/4.2
-LSeg3D [139] 0.0/64.4/0.0 0.9/55.7/0.1 1.8/68.4/0.9
-PLA without caption [85] 39.7/68.3/28.0 24.5/70.0/14.8 25.7/75.6/15.5
-PLA [85] 65.3/68.3/62.4 55.3/69.5/45.9 53.1/76.2/40.8
-WS3D (Ours) 66.1/69.1/65.5 63.3/70.3/58.8 59.3/77.5/51.6
-WS3D-Open (Ours) 68.3/70.6/67.1 64.2/71.4/59.9 59.7/77.9/51.6
-WS3D++ (Ours) 72.3/69.1/73.3 73.3/70.0/63.6 65.2 / 78.9 / 58.8
-Fully-supervised 74.5/68.4/79.1 73.6/72.0/72.8 69.9/75.8/64.9
-Datasets Models Few-shot settings
-B12/N3 B10/N5 B6/N9
-NuScenes 3DGenZ [141] 01.6/53.3/00.8 01.9/44.6/01.0 01.1/52.6/00.5
-3DTZSL [142] 01.2/21.0/0.6 06.4/17.1/03.9 2.61/18.52/03.15
-LSeg-3D [85] 0.6/74.4/0.3 0.0/72.5/0.0 2.66/69.72/0.21
-PLA without caption [85] 25.5/75.8/15.4 10.7/76.0/05.7 8.95/65.83/6.32
-PLA [85] 47.7/73.4/35.4 24.3/73.1/14.5 15.63/60.32/12.38
-WS3D (Ours) 51.3/76.6/30.9 41.5/72.5/20.8 35.2/40.9/29.8
-WS3D-Open (Ours) 55.8/77.5/34.9 47.2/69.3/50.9 39.2/45.3/41.6
-WS3D++ (Ours) 58.6/79.1/38.6 49.7/67.3/53.9 40.6/46.6/51.9
-Fully-supervised 73.7/76.6/71.1 74.8/76.8/72.8 74.6/75.9/72.3
-Figure 12: The WS3D++ corresponding detection and segmentation projections on six rendered 2D views through proposed multi-view rendering.
-It can be demonstrated that our proposed rendering has established an explicit modality association between the final 2D views and 3D views.
-Also, by combining 2D and 3D region proposals, more complete and apparent object-level information can be clearly captured both from 2D views and 3D views.
-Best zoom-in for viewing.
-Figure 13: t-SNE visualization in semantic segmentation of the proposed WS3D under the 5% labeling percentage on S3DIS validation set.
-The diverse feature embeddings are indicated by different colors and are normalized into $[-1,1]$ for better visualization.
-It has been validated that more discriminative features can be acquired for diverse semantic classes with our proposed unsupervised region-aware fine-tuning strategy (demonstrated by WS3D) and our proposed hierarchical vision-language knowledge associated and distilled pre-training (demonstrated by WS3D++).
-TABLE V: Comparison of the performance of instance segmentation, under various levels of supervision on ScanNet validation set.
+ScanNetv2-WS3D++ 69.9 71.8 85.8 78.9 81.7 87.8
+ScanNetv2-WS3D++ (Pr) 75.3 79.8 87.8 85.6 88.8 91.8
+S3DIS WS3D++ 62.8 70.6 75.8 69.7 82.0 86.8
+S3DIS WS3D++ (Pr) 68.7 75.8 78.7 73.8 82.9 87.3
+Semantic KITTI Val. WS3D++ 58.9 65.9 79.3 74.3 76.8 88.8
+Semantic KITTI Val. WS3D++ (Pr) 62.8 72.9 69.1 68.6 83.9 85.5
+Semantic KITTI Test. WS3D++ 66.7 71.8 72.6 71.6 77.9 79.6
+Semantic KITTI Test. WS3D++ (Pr) 70.8 75.9 78.8 79.1 84.3 86.5
+TABLE V: Comparison of current State-of-the-art (SOTA) approaches in the limited reconstruction case for the 3D object detection tasks with different ratios of labeled data.
+The mean average precision (mAP) given also with the mean $\pm$ standard deviation in three runs of diverse random splits are reported.
+Baseline is the circumstance which is trained with merely labelled data.
+It can be demonstrated our proposed pre-training approach have improved the performance of 3D object detection by a considerable margin.
+Datasets Models 5% 10% 20% 100%
+Induct. Transduct. Induct. Transduct. Induct. Transduct. Induct. Transduct.
+SUN RGB-D [111, 110] Baseline 29.9 $\pm$ 1.5 33.5 $\pm$ 0.8 34.4 $\pm$ 1.1 40.7 $\pm$ 0.9 41.1 $\pm$ 0.3 47.5 $\pm$ 0.5 51.7 $\pm$ 0.9 53.8 $\pm$ 0.7
+Merely the Pre-training Stage 35.8 $\pm$ 1.1 38.2 $\pm$ 1.1 39.3 $\pm$ 0.7 45.1 $\pm$ 0.7 46.1 $\pm$ 0.9 56.9 $\pm$ 0.7 51.9 $\pm$ 1.1 65.8 $\pm$ 1.5
+SESS [112] 34.2 $\pm$ 2.0 38.1 $\pm$ 0.7 42.9 $\pm$ 0.8 45.3$\pm$ 0.9 47.9 $\pm$ 0.4 51.6 $\pm$ 0.2 50.7 $\pm$ 0.5 53.7 $\pm$ 0.5
+SESS (Pr) [112] 42.8 $\pm$ 3.2 42.6 $\pm$ 0.7 48.2 $\pm$ 0.8 53.8 $\pm$ 0.9 50.5 $\pm$ 0.6 56.3 $\pm$ 0.7 55.2 $\pm$ 0.3 63.9 $\pm$ 0.6
+3D-IOUMatch [92] 39.1 $\pm$ 1.9 46.3 $\pm$ 0.7 45.5 $\pm$ 1.5 53.5 $\pm$ 0.3 51.6 $\pm$ 0.5 55.6 $\pm$ 0.8 55.6 $\pm$ 0.3 62.3 $\pm$ 0.7
+3D-IOUMatch (Pr) [92] 43.6 $\pm$ 3.2 48.5 $\pm$ 0.9 48.5 $\pm$ 1.8 56.9 $\pm$ 0.5 59.8 $\pm$ 0.7 56.8 $\pm$ 0.8 65.7 $\pm$ 0.5 69.1 $\pm$ 0.7
+SPD [113] 38.5 $\pm$ 0.7 44.3 $\pm$ 0.8 46.0 $\pm$ 1.0 51.7 $\pm$ 0.9 60.9 $\pm$ 0.5 56.9 $\pm$ 0.9 67.9 $\pm$ 0.7 73.9 $\pm$ 0.8
+SPD [113] (Pr) 43.6 $\pm$ 1.2 47.8 $\pm$ 1.7 49.6 $\pm$ 0.8 53.8 $\pm$ 1.1 58.6 $\pm$ 1.6 68.9 $\pm$ 2.5 69.2 $\pm$ 0.9 83.8 $\pm$ 0.5
+WS3D++ (Ours) 52.7 $\pm$ 0.7 54.5 $\pm$ 0.8 53.7 $\pm$ 0.9 63.8 $\pm$ 0.8 69.8 $\pm$ 0.7 72.9 $\pm$ 1.1 68.6 $\pm$ 0.9 82.8 $\pm$ 0.5
+WS3D++ (Ours) (Pr) 56.9 $\pm$ 0.8 59.7 $\pm$ 1.5 58.7 $\pm$ 1.8 68.9 $\pm$ 1.3 68.6 $\pm$ 1.8 73.7 $\pm$ 1.6 72.6 $\pm$ 1.1 83.9 $\pm$ 0.9
+ScanNet-V2 [18] Baseline 27.9 $\pm$ 0.5 30.8 $\pm$ 1.5 31.1 $\pm$ 0.7 33.2 $\pm$ 0.5 41.6 $\pm$ 0.9 44.5 $\pm$ 1.2 45.8 $\pm$ 0.8 48.9 $\pm$ 0.9
+Merely the Pre-training Stage 35.2 $\pm$ 1.5 32.8 $\pm$ 1.1 35.6 $\pm$ 0.8 36.6 $\pm$ 0.6 43.9 $\pm$ 1.1 47.9 $\pm$ 0.8 49.3 $\pm$ 0.7 56.8 $\pm$ 0.9
+SESS [112] 32.2 $\pm$ 0.8 36.8 $\pm$ 1.1 39.7 $\pm$ 0.9 44.5 $\pm$ 0.5 48.2 $\pm$ 0.4 49.6 $\pm$ 0.7 55.8 $\pm$ 0.1 56.8 $\pm$ 0.7
+SESS [112] (Pr) 35.9 $\pm$ 0.9 43.9 $\pm$ 1.2 43.5 $\pm$ 1.2 48.9 $\pm$ 0.5 49.6 $\pm$ 0.6 53.2 $\pm$ 0.7 56.9 $\pm$ 0.6 59.6 $\pm$ 0.9
+3D-IOUMatch [92] 40.0 $\pm$ 0.9 46.3 $\pm$ 0.7 47.2 $\pm$ 0.4 49.6 $\pm$ 0.6 52.8 $\pm$ 1.2 62.9 $\pm$ 0.8 56.9 $\pm$ 1.2 63.7 $\pm$ 0.8
+3D-IOUMatch (Pr) [92] 45.6 $\pm$ 2.0 52.8 $\pm$ 0.8 53.3 $\pm$ 1.5 55.6 $\pm$ 1.7 57.7 $\pm$ 1.2 59.6 $\pm$ 0.8 61.8 $\pm$ 1.6 66.9 $\pm$ 1.3
+SPD [113] 41.5 $\pm$ 0.5 44.3 $\pm$ 0.8 43.2 $\pm$ 1.2 46.2 $\pm$ 0.5 51.9 $\pm$ 0.5 55.6 $\pm$ 0.9 54.8 $\pm$ 0.5 58.6 $\pm$ 0.9
+SPD [113] (Pr) 45.2 $\pm$ 0.7 48.8 $\pm$ 1.9 52.8 $\pm$ 1.6 55.2 $\pm$ 0.7 53.8 $\pm$ 1.8 65.2 $\pm$ 2.5 59.9 $\pm$ 0.5 68.9 $\pm$ 1.6
+WS3D++ (Ours) 55.3 $\pm$ 1.1 57.6 $\pm$ 1.2 59.8 $\pm$ 0.8 66.8 $\pm$ 1.2 64.6 $\pm$ 0.5 72.8 $\pm$ 1.1 68.7 $\pm$ 0.5 76.6 $\pm$ 1.1
+WS3D++ (Ours) (Pr) 66.6 $\pm$ 1.5 68.1 $\pm$ 1.7 67.6 $\pm$ 1.6 74.5 $\pm$ 1.3 69.5 $\pm$ 0.9 76.9 $\pm$ 0.9 73.6 $\pm$ 0.7 77.8 $\pm$ 0.6
+TABLE VI: Comparison of the performance of instance segmentation, under various levels of supervision on ScanNet validation set.
 ’Sup-only-GPC’ denotes the model trained with only labeled data.
@@
 WS3D-Open abandons our feature-aligned contrastive pre-training with the contrastive loss term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [7] feature encoder with the knowledge distillation loss term $\mathcal{L}^{Dist}_{\mathcal{KL}}$.
+WS3D-Open is examined to validate the effect of hierarchical feature-aligned pretraining in improving data efficiency.
+The results of our proposed approach are highlighted in orange.
 Tested Dataset Approaches Ins. Seg. Results with the metric of AP@50%
 1% 5% 10% 15% 20% 30% 35% 40% 100%
-ScanNet Sup-only-GPC [121] 10.8 33.6 42.8 45.3 48.2 49.0 49.5 50.2 56.8
+ScanNet Merely the Pre-training Stage 39.7 46.7 53.8 55.6 58.6 63.5 65.2 65.7 65.8
+Sup-only-GPC–Baseline [103] 10.8 33.6 42.8 45.3 48.2 49.0 49.5 50.2 56.8
+Sup-only-GPC (Pr) [103] 16.8 35.9 33.8 48.9 56.6 57.9 58.8 66.8 68.9
 Mix3D [104] 12.7 34.7 43.1 45.7 48.7 49.6 50.2 51.3 57.6
+Mix3D (Pr) [104] 20.6 39.7 48.9 50.9 53.9 58.9 58.7 55.2 63.9
 GPC [103] 16.9 38.6 44.9 47.2 48.6 49.7 51.2 52.0 57.7
+GPC (Pr) [103] 23.8 45.8 53.8 52.8 54.5 55.1 57.2 57.1 60.6
 SPIB_Ins [114] 17.1 38.9 45.3 47.9 48.9 50.5 51.6 52.7 57.6
-WS3D (Ours) [4] 30.8 45.6 53.5 54.7 53.2 51.9 52.5 53.0 58.7
-WS3D-Open (Ours) 31.2 50.6 55.8 57.9 56.8 58.7 58.9 59.5 65.1
-WS3D++ (Ours) 39.2 54.8 57.3 58 +8.6 +9.7 +9.4 +9.3 +8.3
+SPIB_Ins (Pr) [114] 23.7 52.8 52.9 55.3 54.8 55.8 56.9 58.8 64.8
+WS3D++ (Ours) 45.9 61.8 66.8 67.7 65.1 65.2 62.7 68.7 58.8
+WS3D++ (Ours) (Pr) 48.2 72.2 71.6 70.5 67.9 68.7 71.6 75.8 73.9
 4.4 Data-Efficient 3D Instance Segmentation
-As our method can be integrated seamlessly into various network backbones and applied to different highly-level understanding tasks, we have also integrated our method with Point-Group [90] for the instance segmentation on ScanNet with results shown in Table V.
+As our method can be integrated seamlessly into various network backbones and applied to different highly-level understanding tasks, we have also integrated our method with Point-Group [73] for the instance segmentation on ScanNet with results shown in Table VI.
 Notice that the performance increase is 21.7% when merely 1% data is labeled compared with the sup-only case.
 It further demonstrates that our proposed approaches for the unsupervised branch have effectively exploited the unlabeled data to improve the feature learning capacity of the model.
-As shown in Fig.
-5, our proposed WS3D and WS3D++ both provide explicit boundary guidance for separating diverse kinds of semantic classes, and the instance segmentation performance with very limited labeling percentage is comparable to those fully supervised counterparts.
+Our proposed WS3D and WS3D++ both provide explicit boundary guidance for separating diverse kinds of semantic classes, and the instance segmentation performance with very limited labeling percentage is comparable to those fully supervised counterparts.
 4.5 Data-Efficient 3D Object Detection
-For the data-efficient 3D object detection, following our previous work [99], we evaluate current approaches extensively on SUN RGB-D [137] and ScanNet [20] benchmarks for 3D object detection tasks with the strong 3D object detection backbone VoteNet [126].
+For the data-efficient 3D object detection, following our previous work [81], we extensively evaluate current approaches extensively on SUN RGB-D [111] and ScanNet [18] benchmarks for 3D object detection tasks with the strong 3D object detection backbone VoteNet [115].
 It can be demonstrated the open-vocabulary designs can to some extent boost the 3D object detection performance, which demonstrate the generalization capacity of our fundation model.
-As shown in Table III, we have also tested the ablated approach WS3D-Open which abandons our feature-aligned pre-training term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [9] feature encoder for the knowledge distillation loss term with $\mathcal{L}^{Dist}_{\mathcal{KL}}$.
-It can be demonstrated that the performance degradation can obviously be observed when comparing WS3D-Open to WS3D, which demonstrate the effectiveness of our proposed hierarchical feature aligned pre-training in improve data efficiency.
-TABLE VI: WS3D and WS3D++ ablation studies on ScanNet (Left Value), S3DIS (Middle Value) and Nuscene (Right Value) validation set, for semantic segmentation (Metric: mIoU%) and on ScanNet validation set for the instance segmentation (Metric: AP@50%), both tested with the 5% labeled case.
-Cases Base $w_{i,j}$ $H_{i,j}$ in EF $H_{i,j}$ in UCSL UCSL MS-UCSL SCE WS3D mIoUWS3D++ mIoUWS3D AP@50% WS3D++ AP@50% WS3D mAPWS3D++ mAP%
-No. 1 56.2 / 54.6 / 53.9 62.8 / 58.1 / 56.3 45.6 54.8 45.8 58.9
-No. 2 51.0 / 49.3 / 47.8 57.1 / 52.5 / 50.9 39.9 48.5 40.6 53.6
-No. 3 49.9 / 47.2 / 46.6 53.1 / 50.9 / 47.9 37.0 44.5 38.1 51.7
-No. 4 51.6 / 52.1 / 47.9 56.1 / 54.8 / 47.7 40.1 46.7 41.6 51.5
-No. 5 51.1 / 51.4 / 49.6 57.1 / 53.8 / 49.8 40.7 47.6 40.8 53.6
-No. 6 52.5 / 50.9 / 50.3 56.7 / 53.9 / 49.5 42.2 48.7 39.8 53.8
-No. 7 49.3 / 48.0/ 47.6 55.6 / 53.5/ 50.8 38.1 46.8 39.8 54.3
-No. 8 54.3 / 52.8 / 50.8 60.6 / 55.3 / 52.7 42.9 48.1 41.2 52.7
-No. 9 48.1 / 45.3 / 45.7 50.8 / 47.6 / 50.6 34.8 39.8 33.9 45.3
-TABLE VII: Ablation studies of the designed models of our proposed WS3D++.
-It is demonstrated that the proposed proposal generation approach and local and global contrastive approach can significantly boost final recognition performance.
-Cases 2D Proposal 3D Proposal $\mathcal{L}^{G}_{Ctr}$ $\mathcal{L}^{L}_{Ctr}$ mIoUAP@50% ScanNet B10/N9 Nuscene B6/N9
-No. 1 62.7 / 58.1 / 56.3 58.5 / 55.6 / 53.7 65.2 / 78.9 / 58.8 69.7 / 71.7 / 66.8
-No. 2 55.2 / 51.9 / 50.6 51.3 / 48.8 / 47.9 56.9 / 70.7 / 53.8 66.1 / 66.6 / 64.5
-No. 3 53.6 / 50.6 / 49.6 52.6 / 48.5 / 47.6 52.6 / 48.5 / 47.6 53.6 / 48.5 / 47.6
-No. 5 53.2 / 51.7 / 49.3 50.2 / 46.1 / 45.8 50.8 / 45.8 / 45.5 50.8 / 46.7 / 45.8
-No. 6 51.2 / 49.2 / 49.1 49.2 / 45.8 / 45.2 48.8 / 44.7 / 44.2 48.2 / 43.8 / 43.2
-4.6 Ablation Study
-Ablations: The results for WS3D are summarized in Table VI.
-We have ablated network modules in all combinations of settings as follows.
-Take the ScanNet instance segmentation at AP@50% as examples: Case 1: The full WS3D.
-Case 2: Removing the boundary prediction network, and not using the guidance of $w_{i,j}$.
-The framework still consists of the supervised branch, unsupervised guidance of the energy function based on the predicted confident pseudo label, and contrastive learning.
-This setting leads to a significant drop of 5.7% on AP.
-Case 3: Removing the pairwise term in the energy-based optimization function $E_{sum}$, the AP drops largely by 8.6%.
-Case 4: Removing $H_{i,j}$ in the energy function, the performance drops by 5.5%.
-Case 5: Removing $H_{i,j}$ in the unsupervised contrastive learning, the performance drops by 4.9%.
-Case 6: Conducting contrastive learning only with the region-level feature $\textbf{f}_{g,5}$ and $\textbf{f}_{h,5}$ at the fifth network stage, rather than at multiple stages.
-The performance drops by 3.4%.
-Case 7: Removing the unsupervised region-level contrastive learning branch, the performance drops largely by 7.5%.
-Case 8: Removing the supervised learning branch with the cross-entropy loss, the performance drops by 2.7%.
-The results for WS3D++ are summarized in Table VII.
-Case 9: Only using the supervised branch, the ins.
-seg.
-performance drops significantly by 10.8%.
-Case 10: As demonstrated in Table VII, dropping the 2D or 3D proposals will both result in the performance drops both for the semantic segmentation as well as instance segmentation.
-Case 11: As shown in Table VII, either removing the global contrastive loss or the local contrastive loss results in less effective contrastive representations and affects the final downstream scene understanding performance.
-Analyses: From the above ablations, some important findings are summarized: For WS3D, firstly, not using our designed modules results in a significant performance drop (Cases No.
-3, No.
-7, and No.
-9), which demonstrates the effectiveness of the proposed unsupervised branch and learning strategies to leverage the unlabeled data.
-Secondly, our proposed learning strategies with boundary label $w_{i,j}$ (Case No.
-2), energy function design (No.
-3), high-confidence prediction based energy function design (No.
-4), high-confidence based region-level contrastive learning strategy (No.
-5), multi-stage contrastive learning network design (No.
-6), all have a boost on the overall semantic/instance segmentation performance.
-The results demonstrate that the proposed energy loss is significant for semantic/instance seg.
-performances, because semantic boundary labels are crucial for identifying diverse objects.
-Thirdly, removing the supervision (Case No.
-8), our method still maintains performance with a slight drop of performance by 2.7%.
-It further validates the robustness and feature learning capacity of our approach.
-Last but not least, as demonstrated in Table VII, for our proposed WS3D++, either dropping the 2D/3D proposals, or dropping the global/local contrastive loss result in considerable performance drops.
-It validates the effectiveness of our designs in enhancing discriminative 3D representations.
-4.7 Qualitative and Quantitative Results of the Open-world Recognition Approaches
+As shown in Table V, we have also tested the ablated approach WS3D-Open which abandons our feature-aligned pre-training term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [7] feature encoder for the knowledge distillation loss term with $\mathcal{L}^{Dist}_{\mathcal{KL}}$.
+It can be demonstrated that the performance degradation can be observed when comparing WS3D-Open with WS3D++, which demonstrate the effectiveness of our proposed hierarchical feature aligned pre-training in improving the data efficiency in downstream scene parsing.
+4.6 Qualitative and Quantitative Results of the Open-world 3D Recognition Approaches
 In this Subsection, we further evaluate the performance of the open-world recognition capacity of our proposed approach.
-The results of open-world recognition are shown in Table IV.
-We have compared our work with the previous approach PLA [85] in establishing the sufficient point-language associations for the open-world robot learning.
+The results of open-world recognition are shown in Table I.
+We have also compared our work with the previous approach PLA [69] in establishing the sufficient point-language associations for the open-world robot learning.
 The results demonstrate that our proposed approach has superior performance in open-world recognition.
@@
 It demonstrates that our proposed WS3D++ can enable the scene-level object recognition based on the semantic language queries.
-As further shown in Figure 12, through our effective rendering techniques, which establish the explicit 2D-3D association, the aligned representation of 2D-3D-language co-embeddings can be learned and the objectness can also be enhanced through finding the similarity among diverse views through contrastive learning approaches.
+As further shown in Figure 13, through our effective rendering techniques, which establish the explicit 2D-3D association, the aligned representation of 2D-3D-language co-embeddings can be learned and the object information can also be enhanced through finding the similarity among diverse views through contrastive learning approaches.
 Also, by combining 2D and 3D region proposals, more complete and apparent object-level information can be clearly captured both from 2D views and 3D views.
+It turns out that in the first place, while initializing other various weakly supervised approaches, our proposed approach can realize consistent improvement on the final performance of weakly supervised scene parsing, which demonstrates the superior generalization capacity benefiting from our designed pre-training of our proposed WS3D++ framework.
+In the second pace, the performance of our proposed WS3D++ is comparatively superior compared with the existing weakly supervised comparative approaches listed above.
 4.7 Instance Discrimination Capacity
 We show t-SNE visualizations of the learned latent feature representations for various semantic classes in Fig.
-13.
+12.
 The case study task is the semantic segmentation on the S3DIS dataset with a supervision level of 5%.
 It is demonstrated that more distinctive and better separated point-wise feature embeddings are provided by our proposed unsupervised region-level contrastive learning, which can be attributed to its strong instance discrimination capacity.
-And more separated feature space can be provided with our proposed WS3D++ compared to WS3D and GPC.
+And more separated feature space can be provided and maintained with our proposed WS3D++ compared to WS3D and GPC.
 This strong instance discrimination capacity can be explained by more discriminative feature representations guided by 3D vision-language aligned representations, and is thus more beneficial to high-level semantic and instance segmentation performances both in terms of data efficiency and open-world recognition capacity.
 Also, our proposed hierarichical feature alignment also provides more separated feature space, which means that the feature alignment successfully enhances the final instance discrimination capacity.
+Robot Arm Grasping Example.
+We have deployed our approach for the task of open-world perception in robot grasping in our extended robotic research work.
+We use our proposed approach for segmentation and use the ROS 2 Gazebo-based framework to implement the other components of the system, such as kinematics/dynamics modeling, motion planning, low-level control, point cloud-based pose estimation, etc.
+Our proposed approach has robust performance and decent accuracy in grasping, which demonstrates the potential of our proposed WS3D++ in industrial manipulation applications in grasping and dropping novel objects beyond the training set.
 5 Conclusion
-In this paper, we proposed a general WS3D++ framework for data-efficient 3D scene parsing.
-It involves both the pre-training and the fine-tuning stages.
-In the pre-training stage, we propose the hierarchical feature alignment strategy to acquire accurate regional 3D-linguistic pairs, thus the performance can be enhanced to a large extent.
+In this paper, we proposed a general WS3D++ framework for open-vocabulary and data-efficient 3D scene parsing.
+The whole framework involves both the pre-training and the fine-tuning stages.
+During the pre-training stage, we propose the hierarchical feature alignment strategy to acquire accurate regional 3D-linguistic pairs, thus the performance can be enhanced to a large extent.
 At the same time, we propose an unsupervised boundary-aware energy-based loss and a novel region-level multi-stage semantic contrastive learning strategy, which are complementary to each other to make the network learn more meaningful and discriminative features from the unlabeled data.
 The effectiveness of our approach is verified across three diverse large-scale 3D scene understanding benchmarks under various experiment circumstances.
-Our approach can maximally exploit the unlabeled data to enhance the performance both for 3D point clouds semantic segmentation and instance segmentation, under various labeling percentages in the limited reconstruction case.
+Our approach can maximally exploit the unlabeled data to enhance the performance both for 3D point clouds semantic segmentation and instance segmentation, and object detection under various labeling percentages in the limited reconstruction case.
 Our proposed label-efficient learning framework, termed WS3D++, provides conprehensive baselines for future 3D scene parsing methods when the label is inaccessible or limited.
-Robot Arm Grasping Example
-We deploy our approach for the task of open-world perception in robot grasping.
-We use our proposed approach for segmentation and use the ROS 2 Gazebo-based framework to implement the other components of the system, such as kinematics/dynamics modeling, motion planning, low-level control, point cloud-based pose estimation, etc.
-Our proposed approach has robust performance and decent accuracy in grasping, which demonstrates the potential of our proposed WS3D++ in industrial manipulation applications in grasping and dropping novel objects beyond the training set.
-The video demos are attached.
-Supplementary Material
-In this supplementary material, additional experimental results and details that are not included in the main paper due to space limits are provided as follows:
-  • Details of our experimental datasets (see Section A).
-  • Additional qualitative experimental results (see Section B).
-  • The PFH-based point cloud over-segmentation method to provide the initial over-segmentation results for our proposed WS3D (see Section C).
-  • Details of our data augmentation approach (see Section D).
-  • Training details of the boundary prediction network (see Section E).
-  • The low density issues for the outdoor point clouds captured by LiDAR sensors (see Section F).
-A Experimental Datasets Details
-To demonstrate the effectiveness of our proposed WS3D and WS3D++, we tested it on both indoor and outdoor 3D scene segmentation benchmarks, including S3DIS[21], SemanticKITTI[18], and ScanNet[20] for semantic segmentation, and ScanNet[20] for instance segmentation.
-The detailed information of each dataset is listed as follows:
-Figure 14: Qualitative semantic segmentation results on SemanticKITTI with 5% labels, compared with the fully-supervised arts Cylinder3D [119], and BAAF-Net [89] with semantics indicated by diverse colors.
-The red circles highlight the performance difference between diverse methods.
-ScanNet-V2 is a standard large-scale and widely-acknowledged 3D indoor scene understanding benchmark consisting of more than 1,600 3D scene scans.
-It contains densely reconstructed point clouds from RGB-D images captured by the depth camera.
-For the dataset partition in fully supervised learning, we follow the official partition of ScanNet-V2[20] using 1,201 scans as the training set, 312 scans as the validation set, and 100 scans as the test set.
-SemanticKITTI[18] is a large-scale point cloud understanding benchmark for self-driving applications.
-The point cloud of individual scenes is obtained densely by Velodyne VLP-64 LiDAR.
-The dataset covers long-range road scenes, thanks to the vehicle driving in complex road scenarios with mounted LiDAR to capture the point clouds.
-More than 43,550 labeled LiDAR scans in total are split into 21 sequences.
-Each LiDAR scan contains approximately $10^{5}$ points.
-We follow the official setting[18] that uses the sequences 00 to 07, and 09 to 10 for training, the sequence 08 for validation, and the sequence 11 to 21 for testing.
-S3DIS is a commonly used indoor 3D scene understanding benchmark that contains various indoor scenes.
-It includes six indoor areas which are made up of 271 rooms.
-Each room contains a magnitude of $10^{6}$ points.
-The typical room size is $20\times 15\times 5$ meters.
-In our experiments, we follow GPC[121] using area 5 as the validation set and the other areas as the training set.
-NuScene is a commonly used autonomous driving dataset containing more than 1000 scenes.
-Every scene is elaborately annotated with the 3D bounding boxes for 23 categories and 8 attributes.
-For the base/novel category partition, we overlook the ’otherflat’ class and randomly divide the rest 15 classes into B12/N3 and B10/N5 as in PLA [85].
-Figure 15: Results of our proposed PFH-based over-segmentation in Algorithm 1. Diverse over-segmented regions are indicated by diverse colors.
-B Additional Experimental Results
-We provide the additional experimental qualitative results on ScanNet and SemanticKITTI semantic segmentation and ScanNet instance segmentation.
-In Fig.
-14, the results on SemanticKITTI[18] are shown for the 5% labels case compared with the fully-supervised SOTAs Cylinder3D[119], and BAAF-Net[89].
-From the qualitative results, our method can sometimes offer better segmentation performance than fully-supervised SOTAs for objects such as roads, trees, etc., which demonstrates the effectiveness of our proposed WS3D for outdoor 3D semantic segmentation.
-Comparisons with SOTAs in the fully supervised mode.
-To demonstrate the feature learning capacity of our proposed WS3D, we have also experimented with the 100% labeling percentage for a fair comparison with fully supervised SOTAs.
-The results are shown in Table VIII and the last column of Table I.
-We have fed the whole training set into the supervised branch and unsupervised branch in Fig.
-3 simultaneously.
-Therefore, our proposed energy-based loss and region-level contrastive learning strategy operate as additional optimization guidance for network training.
-Table VIII demonstrates that we can realize at least comparable or even better results, compared with the fully supervised SOTAs.
-Transductive learning.
-Similar to the experimental setting of [121], we have also conducted experiments evaluating the performance of WS3D in transductive learning.
-Different from the inductive learning we tested above that requires the trained model to be generalized to an unseen test set, transductive learning can exploit the testing set when training.
-Compared with inductive learning, we add the test set as part of the unlabeled data in transductive learning.
-As is demonstrated in Table III, the sem.
-seg.
-mIOU becomes higher if the network is learned in a transductive way, both for the fully labeled case with 100% labels and the weakly labeled case with 20% labels.
-It demonstrates that our proposed WSL approaches, including the energy-based boundary-aware loss and the region-level contrastive learning, can leverage the unlabeled data for feature learning effectively in an implicit way to enhance the final segmentation performance.
-Figure 16: In the limited reconstruction case, our proposed approaches WS3D and WS3D++ work well for the low-density SemanticKITTI over-segmentation and boundary prediction (5% label percentage).
-It can also be demonstrated that the explicit boundary information can be clearly captured, and our proposed approach is robust to the density variation of indoor dense and outdoor sparse point clouds.
-C Point Clouds Over-segmentation
-C.1 PFH-based Point Clouds Over-segmentation
-As mentioned in Subsection 3.1 in the main paper, the previous over-segmentation methods depended on the region growing[34], which relied heavily on the accurate normal estimation and were easily influenced by noises.
-In our work, we use the normal, curvature, and point feature histogram[116] simultaneously to provide the initial over-segmentation results for obtaining different regions.
-The point feature histogram (PFH)[116] can be regarded as a robust representation of local geometric features, which is robust to 6D pose transformation and deals well with diverse sampling densities, and neighboring noises.
-We use it as a more robust geometric feature to obtain the over-segmentation result.
-To be more specific, we denote the average normal vector of a point as $\textbf{n}=(n_{x},n_{y},n_{z})$, the PFH feature as $\textbf{f}_{pfh}\in\mathbb{R}^{125\times 1}$.
-The cosine of the angle between the average normal vectors of two adjacent points $P_{i}$ and $P_{j}$ is denoted as $A_{n}$, and the cosine of the angle between two PFH vectors of two adjacent points in the 125-dimensional space is denoted as $A_{pfh}$.
-The Affinity $A(P_{i},P_{j})$ of two adjacent points $P_{i}$ and $P_{j}$ is given as:
-$A(P_{i},P_{j})=\sqrt{A_{n}^{2}+A_{pfh}^{2}}.$ (14)
-According to the Point Cloud Library (PCL)[34], in the process of region growing, points with an affinity larger than a threshold $T_{thres}$ are merged iteratively.
-We set $T_{thres}$ as 1 following the typical setting[34].
-As shown in Algorithm 1, we summarize the detailed procedures for the PFH-based over-segmentation as the following steps:
-Input: The input raw point set $\textbf{P}_{i}=\{p_{i}\},i=1,2,...,N_{i}$. $p_{i}=(x_{i},y_{i},z_{i})$.
-Output: The output pseudo label matrix $\textbf{{L}}_{region}$ for different regions.
-Initialize $P_{seed}=p_{\{seed,i\}},i=1,2,...,N^{seed}$;
-Initialize the pseudo label matrix $\textbf{{L}}_{region}$ as a zero matrix. while not converged do
-       Select $K$ nearest neighbour points $p_{j}$ around the seed points $p_{seed}$ for comparisons based on fast Octree-based $K$ Nearest Neighbour search;
-       for the seed points $p_{j}$ selected do
-             if Condition 1 then
-                  Assign $p_{j}$ the same overseg-label as $p_{seed}$;
-                   if Condition 2 then
-                        Regard the point $p_{j}$ as new seed points;
-            else
-                  Assign $p_{j}$ with a new overseg-label. Regard $p_{j}$ as new seed points;
-            $j\leftarrow j+1$;
-             Update the pseudo label matrix $\textit{{L}}_{region}$.
-return The class label matrix $\textit{{L}}_{region}$ of $\textbf{P}_{i}$ with different clusters.
-Algorithm 1 The Proposed PFH-based Region Growing Over-segmentation to Obtain Regions
-  1. 1.
-    Select the points that have top 2‰ minimum curvatures[34] as the initial seed points $p_{seed}$.
-Utilizing fast $K$ Nearest Neighbour search[120], we obtain the $j_{th}$ neighbouring point $p_{j}$ of the initial seed point $p_{seed}$, and calculate the Affinity $A(p_{seed},p_{j})$ between $p_{seed}$ and $p_{j}$.
-  2. 2.
-    If the Affinity $A(P_{seed},P_{j})$ is larger than $T_{thres}$ (Condition 1), we assign the $j_{th}$ neighbouring point $p_{j}$ with the same label as $p_{seed}$.
-  3. 3.
-    If Condition 1 is satisfied, denote the curvature of $p_{j}$ and $p_{seed}$ as $r_{j}$ and $r_{seed}$, and denote the difference of $r_{j}$ and $r_{seed}$ as $\Delta r=\|r_{j}-r_{seed}\|$.
-If $\Delta r\leq\sigma$ (Condition 2), we set $p_{j}$ as the seed point.
-That means we set the points $p_{j}$ as seed points, only if the difference in curvature is small enough.
-Otherwise, we only assign $p_{j}$ the same label as $p_{seed}$.
-  4. 4.
-    If Condition 1 is not satisfied, we assign the neighbouring point $p_{j}$ with a different label from $p_{seed}$, and also assign $p_{j}$ as the seed point.
-Finally, points are assigned to initial over-segmented regions with labels.
-The simple algorithm is summarized in Algorithm 1.
-The loop will terminate if any one of the following convergence conditions is satisfied, which are proposed:
-  1. 1.
-    All points have been assigned with labels;
-  2. 2.
-    There are no seed points that can be added.
-After the procedures above, we can obtain the label $\textit{{L}}_{region}$, which indicates the separated regions obtained by over-segmentation.
-Some randomly selected over-segmentation results are shown in Fig.
-15.
-The PFH-based over-segmentation can automatically divide the whole point cloud scene into geometrically separated regions as shown in Fig.
-15.
-Different colors indicate different regions.
-TABLE VIII: Comparison of SOTAs methods in the semantic segmentation performance on ScanNet validation set, S3DIS validation set (Area 5), and on SemanticKITTI validation set and test set.
-All results are based on the 100% label ratio.
-Top-two results are highlighted.
-Approaches Venue ScanNet S3DIS SemanticKITTI
-Val. Val. (Area 5) Val. Test
-Minkow-Network [37] CVPR19 72.2 65.4 61.1 63.1
-PointASNL [49] CVPR20 66.4 62.6 - 58.8
-KPConv [144] ICCV19 69.2 67.1 - 46.8
-SPV-NAS [145] ECCV20 - - 64.7 66.4
-Fusion-Net ECCV20 - 67.2 63.7 61.3
-MV-Fusion [44] ECCV20 76.4 65.4 - -
-Cylinder3D [119] CVPR21 - - 65.9 67.8
-BAAF-Net [89] CVPR21 - 72.1 61.2 59.9
-Sup-only-GPC [121] ICCV21 70.9 66.4 65.0 65.4
-GPC [121] ICCV21 74.0 68.8 65.8 67.7
-WS3D ECCV22 76.9 72.9 66.9 69.0
-WS3D++ - 77.6 73.5 67.7 69.8
-D Data Augmentation Details
-D.1 Data Augmentation Approaches
-We perform three kinds of data augmentations to enhance the unlabeled data and to perform region-level contrastive learning, including: 1.
-Geometric rotations and mirroring of the point cloud scene.
-2.
-Adding minor disturbance to the point cloud scene.
-3.
-Random scaling of the point cloud scene.
-[…]
-Adding minor disturbance.
-We add the noise for the normalized point clouds coordinates, which can be represented as $P_{x}^{d}=P_{x}^{m}+T_{d}$.
-$T_{d}$ is the Gaussian noise with the mean of 0 and variance of 0.01.
-We set $T_{d}$ in the range of $[-0.1,0.1]$.
-Random scaling.
-We do random scaling to the point cloud scene.
-Random scaling of ($0.9-1.1$) is done to the point clouds coordinates.
-Denote the scaling parameter as $\eta$, the final augmented point clouds scene $P_{x}^{{}^{\prime}}$ is given as $P_{x}^{{}^{\prime}}=\eta P_{x}^{d}$.
-E Training Details of the Boundary Prediction Network
-[…]
-F Low-Density Issues of Outdoor LiDAR points
-While using the boundary prediction network to extract the semantic boundary for large-scale 3D scenes, the density issue should be taken into consideration because the density variation may influence the quality of the obtained pseudo labels.
-Randomly selected results of SKITTI LiDAR over-segmentation and boundary predictions results are shown in Fig.
-16.
-The density influences over-segmentation performance especially for outdoor low-density point clouds because the 3D point cloud normals, curvatures, and PFH features are all directly related to density.
-For outdoor sparse SemanticKITTI, we have a very simple but effective trick to tackle low-density outdoor LiDAR points: discarding all regions containing less than $N_{thres}$ points, i.e.
-not using these regions in all our network modules.
-($N_{thres}=100$ was set empirically for SKITTI).
-As shown in Fig.
-16, isolated regions with fewer points than $N_{thres}$ are successfully discarded for reliable and robust region-level predictions.
-As shown in Fig.
-16, compared to Karpathy’s approach, our PFH-based over-segmentation better distinguishes similar semantic classes such as road and sidewalk, revealing that our adapted PFH can have a better local feature description capacity than other classical features, both for our proposed WS3D and WS3D++.
-The results of boundary predictions are not much or greatly influenced by density as shown in Fig.
-16.
-Our boundary prediction network learns high-level semantic boundaries, which are not explicitly related to the low-level point density.
-In summary, our proposed WS3D and WS3D++ still work well for the low-density cases.
+The proposed pre-training as well as fine-tuning approach can have a significant boost on the final open-vocabulary and data-efficient semantic scene parsing in term of efficiency, effectiveness, and robustness.